### v1

# KIX: A Metacognitive Generalization Framework

Abstract
Humans and other animals aptly exhibit general intelligence behaviors in solving a variety of tasks with flexibility and ability to adapt to novel situations by reusing and applying high level knowledge acquired over time. But artificial agents are more of a specialist, lacking such generalist behaviors. Artificial agents will require understanding and exploiting critical structured knowledge representations. We present a metacognitive generalization framework, Knowledge-Interaction-eXecution (KIX), and argue that interactions with objects leveraging type space facilitate the learning of transferable interaction concepts and generalization. It is a natural way of integrating knowledge into reinforcement learning and promising to act as an enabler for autonomous and generalist behaviors in artificial intelligence systems.
1 Introduction
Figure 1: Learning for generalist behaviors necessitates building a natural yet abstract understanding of the environment, as well as meta-reasoning and interaction abilities. The metacognitive knowledge-guided approach captures the inherent abstractions and relations between objects to generate interaction goals, enabling the agent to interact with objects and generalize to solve complex tasks.
Flexibility, ability to actively think and evaluate options –learning and reasoning to enable adaptation– are hallmarks of human intelligence. Core to this intelligence is how easily people can find natural decomposition in world environments and tasks, enabling flexible, generalist behaviors. Replicating such autonomy and general intelligence abilities in artificial agents continues to pose significant challenges, which we argue requires a knowledge-centric metacognitive approach.
Meta-reasoning about next actions, action values, and possible next states facilitates flexible and transferable decision making for novel situations, problem solving, concept learning, and knowledge acquisition. This reasoning involves more than just memory lookup. Meta-reasoning is more broadly sense-making in a knowledge system – enabling people to intuitively direct their learning to make sense of relationships between entities, interact with the same entities in different ways, and ground their interactions for both short and long-term goals. Metacognitive reasoning appears to be central to generalization. In contrast, meta-reasoning systems in artificial intelligence are more focused on resource allocation problems. Although resource allocation underlies efficient functioning, we need to ideate metacognitive generalization, figure 1, to enable generalist behaviors in artificial agents. Generalist behaviors have key characteristics which both constitute a basis of devising a metacognitive framework and compartmentalize the behaviors in a useful way.
Our proposal builds on three main principles for solving complex tasks in a flexible manner. Natural decomposition and abstraction: World environments are naturally composed of entities and categorized by their types. It is similar to how a natural hierarchy of objects and classes is formed. Agents interact with entities in the instance space, which inherit properties and relationships from a higher-level type space. Interactions: World environments are essentially knowledge graphs that represent the underlying objects and their relationships. In order to complete its tasks and achieve its goals, an agent’s metacognitive process employs these knowledge graphs to create a plan that specifies which entities it must interact with and what interactions it should have. Composition of interactions: Agents’ flexible behaviors stem from their ability to learn interactions and leverage interaction pathways for graph traversal at the type-space and instance levels as needed for quick adaptation. Integrating these principles into metacognitive learning, reasoning, and control provides levers and connecting points for instilling generalist behaviors in agents.
Our proposed meta-cognitive framework captures the essence of generalist behaviors and may have far-reaching implications across domains. In this article, we provide a description of the key components of the proposed framework, which exemplify generalist characteristics followed by experimental evaluation, as well as how we envision it enabling flexibility through meta-level knowledge guided interactions.
2 Background
In this section, we briefly describe closely related work as well as the general setting of meta-reasoning, graph neural networks, and policy gradient methods.
Related work Compositionality and abstractions are central for generalization learning. Abstraction requires a method for grounding representations to the actual objects. In previous work, Veerapaneni et al. (2020) presented an object-centric model in which object representations were constructed from raw observations and local simpler functions with uniformity, and then used for prediction and planning. Similarly, Vijay et al. (2019) used a symbolic object-centric representation to study the effect of adding prior knowledge through relations in an object graph. These works only consider object space, which inhibits the abstraction of objects and relational structures that occur naturally in the real world and are used for making decisions. Some of the main questions are what an abstraction should entail and how it can be used to compose interactions that can be used as building blocks for leveraging natural relational structures. We attempt to address these questions, which have been largely overlooked in previous works. Moreover, what does natural abstraction mean? It implies preserving relational higher-order structures and abstractions that occur naturally in the real world. Our proposal preserves abstract knowledge in the form of knowledge graphs, where abstractions are grounded using type space. Learning a compositional structure is accomplished by learning over knowledge graphs, with neighborhoods aggregated until stability is reached. Different from previous works, that’s what our proposal entails.
Other methods emphasize aspects which are different from ours, such as encouraging exploration through intrinsic reward or information gain Ten et al. (2022); Oudeyer et al. (2007); Pathak et al. (2017), or motivation Moerland et al. (2023); Zhu et al. (2023); exploring for a variety of experiences by developing internal goals for empowerment Klyubin et al. (2005) ; using a curriculum to progress from easier to more difficult tasks Soviany et al. (2022); Campero et al. (2020) or inducing a learning hierarchy either based on teacher signals Portelas et al. (2020) Pateria et al. (2021); Hutsebaut-Buysse et al. (2022) or rewards Furelos-Blanco et al. (2023). These methods focus on the execution space, influencing reward signals for learning policies in one way or another but largely ignore the higher order structures and relations that objects in the environment have. Our work distinguishes itself by incorporating natural relational structures in the form of knowledge graphs and allowing such meta-level knowledge structures to guide interactions that govern low-level execution.
Meta-reasoning Meta-reasoning is often defined as the set of processes that monitor our thought processes and regulate resource allocation Ackerman and Thompson (2017). Its basic structure Nelson (1990) consists of two levels: meta and object. The meta level’s functions include control and monitoring. The object’s state changes as control passes from the meta to the object level. The monitoring process allows the meta controller to track the status of object-level processes and actions, as well as the consequences of those actions. However, meta-cognitive reasoning in its current form as a resource manager is insufficient to exhibit characteristics of generalist behaviors, necessitating generalist meta-cognition approaches such as the one proposed here.
Graph neural networks A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is defined by a set of nodes $\mathcal{V}$ with node attributes $\mathtt{X}_{v}$ for $v\in\mathcal{V}$ and a set of edges $\mathcal{E}$ between nodes. Edge from node $u\in\mathcal{V}$ to node $v\in\mathcal{V}$ is given by $(u,v)\in\mathcal{E}$ with attributes $\mathtt{e}_{u,v}$. One typical task is to predict the label or value of the entire graph. Graph neural networks (GNN) (Gilmer et al. (2017); Hamilton et al. (2017) and others) use node connectivity as well as node and edge features to learn a permutation-invariant embedding vector for each node $h_{v},v\in\mathcal{V}$ and the entire graph $h_{\mathcal{G}}$. Typically, the representation is learned through message passing or neighborhood feature aggregation, in which node embedding vectors are updated in each layer based on message aggregation from neighbors:
$h_{v}^{(k)}=\mathcal{F}^{\prime}\left(h_{v}^{(k-1)},\bigoplus_{u\in\mathcal{N}(v)}\mathcal{F}^{\prime\prime}\left(h_{v}^{(k-1)},h_{u}^{(k-1)},\mathtt{e}_{u,v}\right)\right)$
where $h_{v}^{(k)}$ is the updated representation of node $v$ embedding in layer $k$, $\mathcal{N}(v)$ is the set of neighboring nodes of node $v$, $h_{u}^{(k-1)}$ is the representation of neighboring node $u$ in layer $k-1$, $\bigoplus$ is a differentiable, permutation invariant aggregation function, $\mathcal{F}^{\prime}$ and $\mathcal{F}^{\prime\prime}$ are differentiable functions usually neural networks. Node features are transformed prior to aggregation, and the methods used to transform and aggregate features differ between graph network variants. Motivated by the attention mechanism Vaswani et al. (2017) in the transformer model, Veličković et al. (2017) employs self-attention over node features and aggregating over neighborhood nodes in graph attention networks (GAT). We use the GATv2 operator Brody et al. (2021), which appears to improve static attention problems in graph attention layers.
Policy gradient methods We consider a standard reinforcement learning setting Sutton and Barto (2018) in which an agent interacts with the environment over a number of steps. A Markov decision process (MDP) is defined as a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma)$ in which $\mathcal{S}$ is a finite state space, $\mathcal{A}$ is a finite discrete action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ is the transition model, $r:\mathcal{S}\times\mathcal{S}\rightarrow\mathbb{R}$ is the reward function and $\gamma\in(0,1)$ is the discount factor. At each step t, the agent takes an action $a_{t}\in\mathcal{A}$ in state $s_{t}\in\mathcal{S}$ according to its policy $\pi$, which is a mapping of states to actions. As a result, the agent reaches state $s_{t+1}\in\mathcal{S}$, receives a scalar reward $r_{t}\in\mathbb{R}$, and accumulates returns; $R_{t}=\sum_{k=0}^{\inf}\gamma^{k}r_{t+k+1}$ represents the accumulated return from step t. The agent’s goal is to maximize the expected return for each state. The state-action value function $Q^{\pi}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is defined as $Q^{\pi}(s,a)=\mathbb{E}[R_{t}|s_{t}=s,a]$, which is the expected return from selecting action $a$ in state $s_{t}$ following policy $\pi$. The value of state $s$ under policy $\pi$ is defined as $V_{\pi}(s)=\mathbb{E}[R_{t}|s_{t}=s]$, which is the expected return from the state $s$. Policy gradient methods, like REINFORCE Williams (1992) , directly learn a parameterized policy $\pi(s;\theta)$ and update its parameters $\theta$ using approximate gradient ascent of a performance measure, $\nabla_{\theta}J(\theta)=\mathbb{E}_{\pi}[G_{t}\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})]$. To reduce the variance in the expected return estimate, a baseline is subtracted from it; typically, a learned value function is used as the baseline. The resulting quantity $A(s,a)=Q(s,a)-V(s)$ can be viewed as an estimate of the advantage of action $a$ in state $s$. This method is called advantage actor critic (A2C) Sutton and Barto (2018); Degris et al. (2012), and the agent acting according to the policy $\pi$ is the actor, while the value baseline is the critic. Several improved methods have been developed to address sample efficiency, stability, and training difficulty. Proximal policy optimization (PPO) Schulman et al. (2017) is an on-policy policy optimization technique that addresses concerns in actor critic methods using clipped objectives. We adapt PPO for learning policies.
3 Model
Our proposal for meta-cognitive generalization centers around the ideas of natural decomposition of tasks that an agent can accomplish through interactions with objects. Meta-policy recommends localized agent-object interactions. The agent executes the interactions following interaction policies and constructs a compositional solution. Here, we discuss the main ideas of the framework (see figure 2) and conjecture that it aptly exhibits task generalization characteristics.
Figure 2: The meta-cognitive generalization framework consists of three levels: meta-level knowledge (K), interaction (I), and execution (X), which are based on levels of abstraction and computation to provide generalization capabilities. The knowledge or type space contains type space knowledge graphs representing higher order abstractions. The instance space contains object-level instance graphs and functions as an intermediary. The execution space is the lowest level abstraction in which the agent acts on meta-level recommendations. Computations proceed through recognition, recommendation, execution, and interactions to gather experiences for updates to interaction and meta policies.
Knowledge An agent can perform low-level actions $u=\{u_{1},u_{2},...\}$ in the environment $\Xi$. The environment consists of objects, which the agent’s perception systems associate with the entity types. Agent maintains its perceived knowledge in the form of knowledge graphs: a graph of objects and their relationships at the instance level, and a graph of entity types and their abstract relationships at the type level. A type space knowledge graph is an ordered pair $G_{K}=(E,\Lambda)$ that consists of a set of entity type nodes, $E=\{E_{1},E_{2},...,E_{m}\}$, and a set of edges or type level relations, $\Lambda=\{\Lambda_{1},\Lambda_{2},...,\Lambda_{\kappa}\}$. The same type of relation between objects and abstract entity types has the same meaning; for example, adjacency relations indicate that two objects are adjacent to each other, whether in instance or type space. Using the knowledge graph and meta-level action of the meta-policy, $\lambda=\{\lambda_{1},\lambda_{2},...,\lambda_{p}\}$, the meta-level of the agent determines how it will interact with an underlying object. To interact with an object, the agent activates an interaction type, $\lambda$, following the meta-policy guidance. For an interaction, the agent will perform a series of low-level actions. For example, if there is a door, the agent can activate ‘open the door’ at the abstract level, even though it requires many low-level actions such as moving and toggling. The agent uses an activation relation to keep track of which objects it wants to interact with and what type of interaction. Thus, the agent’s knowledge space $\mathcal{K}=(A,E,\Lambda,\lambda,G_{K})$ comprises of a type space knowledge graph, a set of entity types, abstract relations, and interactions.
Interaction The environment consists of objects, $e=\{e_{1},e_{2},...,e_{n}\}$, their types, $\text{type}(e_{i})\in E,\text{for}\ e_{i}\in e$, and is partially observable. Agent keeps track of the visible objects in its instance space and builds an instance graph, which is an ordered pair of nodes (observed objects) and edges (observed object relations), $G_{I}=(e,\Lambda_{I})$. Thus, the agent’s instance space, $\mathcal{I}=(A,e,\Lambda_{I},\lambda,G_{I})$, is made up of its instance space knowledge graph, relationships between objects, and available interactions. After constructing the knowledge graphs, the agent’s goal is to identify the objects for interaction which is done by the meta-level. Meta-level recommends an interaction goal for the agent. For the recommended interaction goals, the agent learns interaction policies for interacting with objects, as well as the relationships between interaction types and entity types. With the ability to interact with objects, the agent accumulates interaction knowledge to update the meta-level policy. Therefore, when viewing tasks as meta-policy guided interactions with objects, problems are naturally decomposed into interaction goals.
eXecution By pursuing interaction goals as directed by the meta-level, the agent learns to interact with objects and act at the execution level. The execution space for the agent is its partially observed view $X$. During an interaction, the agent takes actions $u$ on low-level states $s$. After learning both the interaction and meta policies, the agent acts in the environment, aided by interaction with objects. Therefore, unlike the models that are tightly coupled with low-level representations through end-to-end policies, agent actions are aided by meta-level knowledge-guided object interactions, providing a mechanism for generalist behaviors.
Mapping between execution space and instance space Assume that the agent has an object recognition system that can access its partially observed view $X$. The agent generates an instance-level graph using the observed objects. The function $\Phi:X\to G_{I}$ maps the agent’s object recognition to its instance graph by capturing observed objects and their relationships. As the agent interacts with objects, the environment changes, and its recognition system captures a new partially observed view. The mapping function then generates the corresponding instance graph.
Mapping between instance space and type space The function $\Psi:G_{I}\to G_{K}$ provides a mapping between instance and type graphs. If the agent encounters a novel object for which it does not have a corresponding typed entity, it must either create a new type in its type space or assign the object to a type that is closely related. In the framework, $\Psi$ is learnable and represents a belief of mapping objects to their types. However, in this article, we treat the mapping as a fixed mapping from object nodes in $G_{I}$ to their corresponding type nodes in $G_{K}$.
Proposal for activation object and interaction Using the meta-policy, the agent’s recommendation function $\Upsilon:G_{K},G_{I},\pi\to e_{i},\lambda_{p}$ recommends a candidate object $e_{i}$ for interaction and interaction goal $\lambda_{p}$. The recommender iterates through all visible objects in the agent’s instance graph $G_{I}$, then activates the object by adding an activation relationship between the agent and the object to the instance graph, as well as a corresponding activation relation between the agent and the object’s type node in the type space graph $G_{K}$. Activated $G_{K}$ represents the activated candidate object and serves as the input state for meta-policy, which returns both the state value and meta-action probabilities. The state value represents the value of activating the respective object. Thus, the recommender proposes an object for interaction $e_{i}$ by sampling from the values of activating objects, and the proposed interaction $\lambda_{p}$ is sampled from the meta-action probabilities of the proposed object. Acting on these recommendations, the agent collects interaction experiences $G_{K},\lambda$, that is, interaction trajectories $G_{K}^{0}\curvearrowright G_{K}^{1}...\curvearrowright G_{K}^{t}$, and learns meta-level policy.
Interactions At each step, the agent activates the object $e_{i}$ for interaction $\lambda_{p}$, which is its interaction goal, $g^{i}_{p}=(e_{i},\lambda_{p})$. Following this interaction goal, the agent interacts with the environment. After each round of interaction, the status of the interaction is passed to the meta level. Using the meta-policy, the meta-level recommender suggests next target object and interaction. The process continues until the task is completed or the environment is terminated. Throughout the process, the agent collects experience trajectories at both the interaction and meta levels. The policies are then updated based on the collected experiences.
Meta-policy learning Meta-level representations of type space graphs with activation relations are learned using a message-passing graph neural network. Nodes have abstract level attributes such as type encoding, while edges have edge type encoding for different types of relationships. We use GAT convolutional layers Brody et al. (2021); Fey and Lenssen (2019) to learn graph representations given as,
$h_{v}^{\prime}=\alpha_{v,v}\theta h_{v}+\sum_{u\in\mathcal{N}(v)}\alpha_{u,v}\theta h_{u}$
where $h_{v}^{\prime}$ is the updated embedded representation of node $v$, $\mathcal{N}(v)$ is the set of neighbors of node $v$, and $h_{u}$ is the representation of the neighboring node. $\theta$ are learnable weights shared by source and target nodes in the neighborhood, and the attention coefficients $\alpha_{u,v}$ are computed using weighted node features.
The hidden representation after the final graph attention layer is pooled for graph rollout, which is then passed to the fully connected actor and critic layers of the model. The meta policy $\pi_{\theta}(.|G_{K})$, parameterized by $\theta$, returns a state value and an interaction or meta-action, $\lambda$, distribution. The recommender selects a target object and an interaction $\lambda_{p}$, and the agent works through low-level interactions based on guidance. At the end of the interaction, the status of the interaction is returned to the meta level, and the agent collects the trajectory of meta-level experiences, $\mathcal{D}_{\tau_{m}}$. An estimate of advantages is computed as $A_{m}=r_{m}+\gamma_{m}V_{\pi_{\theta}}(G_{K}^{\prime})-V_{\pi_{\theta}}(G_{K})$.
Meta level PPO optimizes a surrogate loss:
$\mathcal{L}_{m}^{\text{actor}}(\theta)=-\mathbb{E}_{\mathcal{D}_{\tau_{m}}}\left[\min\left(r_{\theta}A_{m},\text{clip}(r_{\theta},1-\epsilon,1+\epsilon)A_{m}\right)\right]$
with $r_{\theta}=\frac{\pi_{\theta}(.|G_{K})}{\pi_{\theta}^{r}(.|G_{K})}$ being the ratio of probabilities of actions at meta level roll-out. The critic loss is the mean squared error between the estimated values and the returns:
$\mathcal{L}_{m}^{\text{critic}}(\theta)=\mathbb{E}_{\mathcal{D}_{\tau_{m}}}\left[\frac{1}{2}(V_{\pi_{\theta}}(G_{K})-R_{m})^{2}\right]$
Considering the mean entropy $\mathcal{H}_{\theta}$, the overall minimization objective at the meta-level becomes:
$\mathcal{L}_{m}(\theta)=\mathcal{L}_{m}^{\text{actor}}(\theta)-\zeta_{m}^{h}\mathcal{H}_{\theta}+\zeta_{m}^{c}\mathcal{L}_{m}^{\text{critic}}(\theta)$
where $\zeta_{m}^{h}$, $\zeta_{m}^{c}$ are the entropy and value coefficients, respectively. The model parameters $\theta$ are optimized as $\theta\leftarrow\theta+\alpha_{m}\nabla_{\theta}L_{m}(\theta)$, and the updated meta-policy is used for the next episode of learning.
Interaction policy learning Interaction policies are a repository of policies for different interactions which can be recommended by the meta-level, $\mu_{\phi}=\{\mu_{\phi_{1}},\mu_{\phi_{2}},...,\mu_{\phi_{p}}\}$ and each policy is parameterized separately. When an interaction goal with an object $e_{i}$ and interaction $\lambda_{p}$ is recommended, $g^{i}_{p}=(e_{i},\lambda_{p})$, the agent attempts to achieve it by executing a series of low-level actions from $u$. The agent observes a low-level state $s$ from a convolutional neural network embedding, samples an action $u\sim\mu_{\phi_{p}}(.|s)$ using the respective interaction policy, acts in the environment, and collects the trajectory of experiences $\mathcal{D}_{\tau_{p}}$. The experiences collected are used to learn interaction policies. An estimate of advantages is computed as $A_{p}=r_{p}+\gamma_{p}V_{\mu_{\phi_{p}}}(s^{\prime})-V_{\mu_{\phi_{p}}}(s)$.
Each interaction level PPO optimizes a surrogate loss:
$\mathcal{L}_{p}^{\text{actor}}(\phi_{p})=-\mathbb{E}_{\mathcal{D}_{\tau_{p}}}\left[\min\left(r_{\phi_{p}}A_{p},\text{clip}(r_{\phi_{p}},1-\epsilon,1+\epsilon)A_{p}\right)\right]$
with $r_{\phi_{p}}=\frac{\mu_{\phi_{p}}(.|s)}{\mu_{\phi_{p}}^{r}(.|s)}$ is the ratio of probabilities of actions of policy and roll-out policy. The critic loss is the mean squared error between the estimated values and the returns:
$\mathcal{L}_{p}^{\text{critic}}(\phi_{p})=\mathbb{E}_{\mathcal{D}_{\tau_{p}}}\left[\frac{1}{2}(V_{\mu_{\phi_{p}}}(s)-R_{p})^{2}\right]$
Considering the mean entropy $\mathcal{H}_{\phi_{p}}$, the overall minimization objective becomes:
$\mathcal{L}_{p}(\phi_{p})=\mathcal{L}_{p}^{\text{actor}}(\phi_{p})-\zeta_{p}^{h}\mathcal{H}_{\phi_{p}}+\zeta_{p}^{c}\mathcal{L}_{p}^{\text{critic}}(\phi_{p})$
for each interaction $p$, where $\zeta_{p}^{h}$, $\zeta_{p}^{c}$ are the entropy and value coefficients, respectively. For each interaction policy, the model parameters $\phi_{p}$ are optimized as $\phi_{p}\leftarrow\phi_{p}+\alpha\nabla_{\phi_{p}}L_{p}(\phi_{p})$, and the updated interaction policies are used for the next episode of learning.
As a result, the agent follows a trajectory of activated interactions with objects and aids the formation of a task solution through meta-guided interactions. The proposed framework is generic to be applied to a wide range of domains.
4 Experiments
Experimental set up To assess the capabilities of the proposed agents, we require a challenging environment with ingredients for higher order reasoning to solve tasks. Our goal here is to investigate whether the introduction of meta-knowledge guided interactions benefits agents. MiniGrid world environments Chevalier-Boisvert et al. (2018) are suitable for our purpose. These are procedurally generated environments in which the agent interacts with objects such as doors, keys, balls, and boxes.
Figure 3: MiniGrid ObstructedMaze Full environment is an appropriate testbed because it is partially observable, requires complex reasoning, and new tasks can be created.
The environments are challenging due to partially observability, sparsity of rewards and random goal locations. The MiniGrid-ObstructedMaze-Full environment, figure 3, in particular, is a difficult environment with only a few claims for solving it. Due to the large state space size, stochasticity and partial observability, agents cannot memorize but must learn.
In the environment, the agent is initially placed in the middle room, with a 7x7 view. At the low level, the agent can move forward, left, right, pick, drop, and toggle, and its goal is to pick up the blue ball, which is randomly placed in one of the corner rooms. Some of the doors are locked and require a key of the same color to open. Keys are hidden inside boxes, but the agent cannot see inside the boxes without toggling. The original task is referred as Task 0. Furthermore, we identify the following additional task scenarios that were not seen during training but were presented to the agent during testing:
  • No context in training (Task 1) - a relation that the agent did not come across during training. It is simulated by placing the goal inside an empty box. During training, the blue ball was never in a box, although the agent encountered boxes and balls during training, but boxes contained keys.
  • Different context in testing (Task 2) - using a relation that the agent encountered during training but in a new context during testing. It is simulated by locking doors in the central room and placing keys directly in the room. During training, doors in the central room were never locked, nor was there direct access to keys.
  • Dynamic changes in the environment (Task 3) - when some event occurs, the environment changes. During training, the goal location remained unchanged. To simulate, as soon as the agent enters the room with the goal, relocate the goal to one of the neighboring rooms.
Agent training
Initialize model weights $\theta,\phi$ , parameters, $T$ for  $t=1$ to $T$ do
       // collect exp for meta exp size K
       for $k=1$ to $K$ do
             make instance graph $G_{I}\leftarrow\Phi(X)$ map $G_{I}$ to $G_{K}$, $G_{K}\leftarrow\Psi(G_{I})$ proposal $(e_{i},\lambda_{p})\leftarrow\Upsilon(G_{K},G_{I},\pi)$ // 1st variant
             interaction exp following $\mu_{\lambda_{p}}$ // 2nd variant
             reaching exp following $\mu_{\text{reach}}$ if success then
                   interaction exp following $\mu_{\lambda_{p}}$
             end if
            collect exp $\mathcal{D}_{\tau_{p}}$ and compute $A_{p}$
       end for
      collect meta exp $\mathcal{D}_{\tau_{m}}$ and compute $A_{m}$ // Update policies, both variants
       update $\pi$ by optimizing $\mathcal{L}_{m}(\theta)$ for  $\lambda_{p}$ do
             update $\mu_{\lambda_{p}}$ by optimizing $\mathcal{L}_{p}(\phi_{p})$
       end for
      // 2nd variant only
       update $\mu_{\text{reach}}$ by optimizing $\mathcal{L}_{\text{reach}}(\phi_{\text{reach}})$
end for
Algorithm 1 KIX
We devise two variants, KIX1 and KIX2. Both variants learn meta-policy and interaction policies. However the second variant additionally learns a reachability policy, so it interacts with an object once it has reached it. All meta, interactions, and reachability policies are implemented as actor-critic PPO policies. Algorithm 1 provides a high-level overview of the implementation. The input state embeddings for interaction and reachability actor-critic networks are constructed by a convolutional network with three Conv2D layers with 16, 32, and 64 output channels, a max pooling layer added after the first convolution layer, and exponential linear units used as activation functions. The meta-level recommends various objects for interaction, so an activation indicator for the object of interest is added to a new input channel, creating a four-channel input. Actors and critics are fully connected networks, each with one hidden layer. Meta-level policies are learned over type space knowledge graphs. In a type space graph, node object type information is encoded as an attribute vector. Edges in type space graph represent type-level relationships between nodes, such as "visible," "adjacent," "carrying," and "activated." A one-hot encoded edge attribute vector is used to encode these relationships. Embeddings for meta-level actor-critic are provided by two graph attention convolution layers, each with 16-dimensional embeddings and four attention heads. During message passing, these layers share weights between source and target. The final graph attention convolution layer is used for graph readout with a global add pool. Meta-level actors and critics are fully connected networks, with the critic learning a single value and the actor learning the probability of taking a meta-level action, such as ’pickup’, ’drop’, ’reveal’, ’open’, or ’open with key’. The meta-level actions serve as interaction goals. The environment rewards the agent with $r=1-(c/c_{T})$ where $c$ is the number of steps taken and $c_{T}$ is the max possible steps. For interaction policy learning, the agent receives a reward of 1 when it has a successful interaction. At the meta-level it is rewarded with 0.1 for a successful proposed interaction and $1+r$ for achieving the environment goal. A base agent is also implemented, with a similar network and a PPO policy, and it receives environmental rewards. Grid search is used to select hyperparameters. All the agents are trained for the same number of environment steps using 32 parallel workers.
Figure 4: The return profiles of agents performing various evaluation tasks show that the proposed agents have higher returns on all tasks and benefit from meta-guided interactions.
Meta-guided interactions aid agents Our proposal is motivated by the idea that, rather than acting directly at the execution level, meta-level guided interactions with objects aid agents in generalizing while performing tasks. Because the agent’s environmental return is determined by the number of steps taken to complete the task, completing tasks in fewer steps yields a higher return. Therefore, the distribution of environmental returns is a reliable indicator of whether meta-level interaction recommendations benefit agents. We evaluate the distribution of environmental returns for each agent across all four tasks. For each task, the trained agents are run for 12K episodes by parallel workers, and their environmental rewards are recorded. In order to understand the effects of meta-guided interactions and to ensure a fair comparison between agents where differences in success rates are not a factor, we examine the top $k$ returns with $k=100$. Figure 4 illustrates the return profiles of agents. It is clear from the distribution of returns that the proposed agents have received a higher return on all tasks than the base agent, furthermore, this is even more so for dynamic tasks like task 3. It demonstrates that meta-guided interactions not only help agents complete their task goals in fewer steps, but also aid them to perform a wider range of tasks in static or dynamic environments. Furthermore, the higher returns of the second variant demonstrate that agents benefit from the introduction of reachability as a separate interaction. In the absence of a separate reachability policy, reaching is implicit in each interaction policy; however, separating allows for a shared learning of reachability followed by interactions which could be beneficial particularly in dynamically changing environments.
Figure 5: Wasserstein distances between traversal distributions, when agents complete tasks 1-3, and the trained task 0, provide comparative insights.
Agents profile While performing the tasks, agents interact with objects in the environment according to the recommended interaction policies; as a result, agents navigate and visit rooms. Because interactions are recommended by the meta-level, which is only trained on the training task, tracking agents’ room visits across all tasks provides an insight about their learning. Let $p_{0}$ be the distribution of agent room visits for trained task 0, and $p_{i},i\in\{1,2,3\}$ be the distributions for tasks 1–3. For both variants of the agent, we then compute the Wasserstein distance, a metric from optimal transport theory, between the visit distribution for tasks 1-3 and the trained task 0, as shown in figure 5. The distribution of visits to each agent is expected to be different from the training task, due to changes in the evaluation tasks. However, our interest is in how differently the two variants behave in evaluation tasks, as indicated by the computed metric. In task 1, the goal room remained unchanged, but a new relation was added to the evaluation. For both variants, the computed metric value for task 1 is similar, indicating that they transfer their learning to this task in a similar manner. However, there is a noticeable difference for tasks 2 and 3. In task 2, doors in the central room are locked with keys kept in the room - a static task, whereas in task 3, the goal is dynamically relocated based on observance - a dynamic task. The computed metric for the dynamic task for the second variant is smaller than the first variant, implying that its evaluation visit distribution is closer to that of the training task. It suggests that the learning in the second variant can better capture dynamic changes. When combined with the return profiles, it reinforces that having an additional reachability policy benefits the agent in a wider range of tasks, particularly dynamic tasks, and improves generalization.
5 Conclusion
In this work, we proposed a metacognitive framework that promotes generalist behaviors through three main contributions. First, generalist and flexible behaviors necessitate abstraction and composition, but current models lack natural abstractions and compositions that can preserve both object abstractions and relational knowledge structures. In our model, natural abstractions and compositions are represented and grounded using type space knowledge graphs which also preserve relational knowledge between objects. Second, rather than directly acting in the raw environment, interactions with objects play an important role in how an agent navigates the environment to complete tasks. To accomplish this, we propose an intermediary interaction level. The meta-level of the agent learns from type space knowledge graphs and recommends what objects to interact with, as well as what interaction should be. In this sense, our model also functions as an interaction goals generating machine. Third, interaction policies learned through interaction goals serve as a repository of policies for the agent to use in the execution space, providing for the necessary loose coupling for generalization via interactions as intermediaries. The agent’s meta-level recommends interactions based on knowledge structures and learned meta-policy, and the execution level applies interaction policies to complete tasks. We demonstrated that metacognitive knowledge guided interactions aid agents in generalization. Our framework is modular and adaptable, with the potential to have implications across domains through the induction or integration of knowledge structures, whether extrinsic, intrinsic, or derived, as well as control variables at the meta or interaction level.

### v2

@@
-KIX: A Metacognitive Generalization Framework
+KIX: A Knowledge and Interaction-Centric Metacognitive Framework for Task Generalization
 Abstract
-Humans and other animals aptly exhibit general intelligence behaviors in solving a variety of tasks with flexibility and ability to adapt to novel situations by reusing and applying high level knowledge acquired over time.
-But artificial agents are more of a specialist, lacking such generalist behaviors.
+People aptly exhibit general intelligence behaviors in solving a variety of tasks with flexibility and ability to adapt to novel situations by reusing and applying high-level knowledge acquired over time.
+But artificial agents are more like specialists, lacking such generalist behaviors.
 Artificial agents will require understanding and exploiting critical structured knowledge representations.
 We present a metacognitive generalization framework, Knowledge-Interaction-eXecution (KIX), and argue that interactions with objects leveraging type space facilitate the learning of transferable interaction concepts and generalization.
-It is a natural way of integrating knowledge into reinforcement learning and promising to act as an enabler for autonomous and generalist behaviors in artificial intelligence systems.
+It is a natural way of integrating knowledge into reinforcement learning and is promising to act as an enabler for autonomous and generalist behaviors in artificial intelligence systems.
 1 Introduction
-Figure 1: Learning for generalist behaviors necessitates building a natural yet abstract understanding of the environment, as well as meta-reasoning and interaction abilities.
-The metacognitive knowledge-guided approach captures the inherent abstractions and relations between objects to generate interaction goals, enabling the agent to interact with objects and generalize to solve complex tasks.
+Figure 1: Learning generalist behaviors requires developing a natural and abstract understanding of the environment, as well as meta-reasoning and interaction abilities.
+The metacognitive knowledge-guided approach exploits inherent abstractions and relationships between objects to generate interaction goals, enabling the agent to interact with objects and generalize to solve complex tasks.
 Flexibility, ability to actively think and evaluate options –learning and reasoning to enable adaptation– are hallmarks of human intelligence.
-Core to this intelligence is how easily people can find natural decomposition in world environments and tasks, enabling flexible, generalist behaviors.
-Replicating such autonomy and general intelligence abilities in artificial agents continues to pose significant challenges, which we argue requires a knowledge-centric metacognitive approach.
-Meta-reasoning about next actions, action values, and possible next states facilitates flexible and transferable decision making for novel situations, problem solving, concept learning, and knowledge acquisition.
-This reasoning involves more than just memory lookup.
-Meta-reasoning is more broadly sense-making in a knowledge system – enabling people to intuitively direct their learning to make sense of relationships between entities, interact with the same entities in different ways, and ground their interactions for both short and long-term goals.
+Core to this intelligence is how easily people can find natural decomposition in world environments and tasks; this enables flexible, generalist behaviors.
+Replicating such autonomy and general intelligence abilities in artificial agents continues to pose significant challenges, which we argue requires a knowledge and interaction-centric metacognitive approach.
+Meta-reasoning about next actions, their values, and possible next states facilitates flexible decision-making in novel situations, problem solving, concept learning, and knowledge acquisition.
+This reasoning extends beyond mere memory lookup.
+Meta-reasoning is more broadly sense-making in a knowledge system - enabling people to intuitively direct their learning to make sense of relationships between entities, interact with the same entities in diverse ways, and ground their interactions for both short- and long-term goals.
 Metacognitive reasoning appears to be central to generalization.
-In contrast, meta-reasoning systems in artificial intelligence are more focused on resource allocation problems.
-Although resource allocation underlies efficient functioning, we need to ideate metacognitive generalization, figure 1, to enable generalist behaviors in artificial agents.
-Generalist behaviors have key characteristics which both constitute a basis of devising a metacognitive framework and compartmentalize the behaviors in a useful way.
-Our proposal builds on three main principles for solving complex tasks in a flexible manner.
-Natural decomposition and abstraction: World environments are naturally composed of entities and categorized by their types.
-It is similar to how a natural hierarchy of objects and classes is formed.
+In contrast, meta-reasoning systems in artificial intelligence are primarily focused on resource allocation problems.
+While resource allocation underpins efficient functioning, it is crucial to ideate metacognitive generalization, Figure 1, to enable generalist behaviors in artificial agents.
+Our proposal builds on three main principles for solving complex tasks flexibly.
+Natural Decomposition and Abstraction: World environments are naturally composed of entities categorized by their types, similar to a natural hierarchy of objects and classes.
 Agents interact with entities in the instance space, which inherit properties and relationships from a higher-level type space.
-Interactions: World environments are essentially knowledge graphs that represent the underlying objects and their relationships.
-In order to complete its tasks and achieve its goals, an agent’s metacognitive process employs these knowledge graphs to create a plan that specifies which entities it must interact with and what interactions it should have.
-Composition of interactions: Agents’ flexible behaviors stem from their ability to learn interactions and leverage interaction pathways for graph traversal at the type-space and instance levels as needed for quick adaptation.
-Integrating these principles into metacognitive learning, reasoning, and control provides levers and connecting points for instilling generalist behaviors in agents.
-Our proposed meta-cognitive framework captures the essence of generalist behaviors and may have far-reaching implications across domains.
-In this article, we provide a description of the key components of the proposed framework, which exemplify generalist characteristics followed by experimental evaluation, as well as how we envision it enabling flexibility through meta-level knowledge guided interactions.
+Interactions: World environments are essentially knowledge graphs representing underlying objects and their relationships.
+An agent’s metacognitive process employs these knowledge graphs to create a plan specifying which entities to interact with and the nature of these interactions to complete tasks and achieve goals.
+Composition of Interactions: Agents’ flexible behaviors stem from their ability to learn interactions and leverage interaction pathways for graph traversal at both the type level and the instance level, enabling quick adaptation.
+Integrating these principles into metacognitive learning, reasoning, and control provides the levers and connecting points for instilling generalist behaviors in agents.
+The proposed metacognitive framework captures the essence of generalist behaviors and may have far-reaching implications across various domains.
+In this article, we describe the key components of the framework that exemplify generalist characteristics, followed by an experimental evaluation as well as how we envision it enabling flexibility through meta-level knowledge guided interactions.
 2 Background
-In this section, we briefly describe closely related work as well as the general setting of meta-reasoning, graph neural networks, and policy gradient methods.
-Related work Compositionality and abstractions are central for generalization learning.
-Abstraction requires a method for grounding representations to the actual objects.
+In this section, we briefly describe closely related work and the general setting of meta-reasoning, graph neural networks, and policy gradient methods.
+Related work
+Compositionality and abstractions are central to generalization learning, with abstraction requiring a mechanism for grounding representations in actual objects.
 In previous work, Veerapaneni et al.
-(2020) presented an object-centric model in which object representations were constructed from raw observations and local simpler functions with uniformity, and then used for prediction and planning.
+(2020) presented an object-centric model where object representations were constructed from raw observations and simpler local functions with uniformity, which were then used for prediction and planning.
 Similarly, Vijay et al.
-(2019) used a symbolic object-centric representation to study the effect of adding prior knowledge through relations in an object graph.
-These works only consider object space, which inhibits the abstraction of objects and relational structures that occur naturally in the real world and are used for making decisions.
-Some of the main questions are what an abstraction should entail and how it can be used to compose interactions that can be used as building blocks for leveraging natural relational structures.
-We attempt to address these questions, which have been largely overlooked in previous works.
-Moreover, what does natural abstraction mean?
-It implies preserving relational higher-order structures and abstractions that occur naturally in the real world.
-Our proposal preserves abstract knowledge in the form of knowledge graphs, where abstractions are grounded using type space.
-Learning a compositional structure is accomplished by learning over knowledge graphs, with neighborhoods aggregated until stability is reached.
-Different from previous works, that’s what our proposal entails.
-Other methods emphasize aspects which are different from ours, such as encouraging exploration through intrinsic reward or information gain Ten et al.
+(2019) used a symbolic object-centric representation to study the impact of adding prior knowledge through relations in an object graph.
+These studies focus primarily on object space, which inhibits the abstraction of objects and relational structures naturally occurring in the real world and used for decision-making.
+Key questions include what a natural abstraction should be and how it can be used to compose interactions as building blocks leveraging natural relational structures.
+We address these largely overlooked questions by preserving abstract knowledge and relational higher-order structures in the form of knowledge graphs, where abstractions are grounded using type space.
+Learning a compositional structure is achieved through learning over knowledge graphs, with neighborhood aggregation until stability is reached.
+This distinguishes our approach from previous works.
+Other methods emphasize different aspects, such as encouraging exploration through intrinsic reward or information gain Ten et al.
 (2022); Oudeyer et al.
 (2007); Pathak et al.
-(2017), or motivation Moerland et al.
+(2017), motivation Moerland et al.
 (2023); Zhu et al.
-(2023); exploring for a variety of experiences by developing internal goals for empowerment Klyubin et al.
-(2005) ; using a curriculum to progress from easier to more difficult tasks Soviany et al.
+(2023), developing internal goals for empowerment Klyubin et al.
+(2005), using a curriculum to progress from easier to more difficult tasks Soviany et al.
 (2022); Campero et al.
-(2020) or inducing a learning hierarchy either based on teacher signals Portelas et al.
+(2020), or inducing a learning hierarchy based on teacher signals Portelas et al.
 (2020) Pateria et al.
@@
 (2023).
-These methods focus on the execution space, influencing reward signals for learning policies in one way or another but largely ignore the higher order structures and relations that objects in the environment have.
-Our work distinguishes itself by incorporating natural relational structures in the form of knowledge graphs and allowing such meta-level knowledge structures to guide interactions that govern low-level execution.
-Meta-reasoning Meta-reasoning is often defined as the set of processes that monitor our thought processes and regulate resource allocation Ackerman and Thompson (2017).
+These methods focus on the execution space, influencing reward signals for learning policies, but largely ignore the higher-order structures and relations between objects in the environment.
+Our work distinguishes itself by incorporating natural relational structures through knowledge graphs and leverages meta-level knowledge structures to guide interactions that govern low-level execution.
+Meta-reasoning
+Meta-reasoning is often defined as the set of processes that monitor our thought processes and regulate resource allocation Ackerman and Thompson (2017).
 Its basic structure Nelson (1990) consists of two levels: meta and object.
@@
 The monitoring process allows the meta controller to track the status of object-level processes and actions, as well as the consequences of those actions.
-However, meta-cognitive reasoning in its current form as a resource manager is insufficient to exhibit characteristics of generalist behaviors, necessitating generalist meta-cognition approaches such as the one proposed here.
-Graph neural networks A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is defined by a set of nodes $\mathcal{V}$ with node attributes $\mathtt{X}_{v}$ for $v\in\mathcal{V}$ and a set of edges $\mathcal{E}$ between nodes.
+However, current meta-cognitive reasoning, primarily as a resource manager, is insufficient for exhibiting generalist behaviors, necessitating generalist meta-cognition approaches such as the one proposed here.
+Graph neural networks
+A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is defined by a set of nodes $\mathcal{V}$ with node attributes $\mathtt{X}_{v}$ for $v\in\mathcal{V}$ and a set of edges $\mathcal{E}$ between nodes.
 Edge from node $u\in\mathcal{V}$ to node $v\in\mathcal{V}$ is given by $(u,v)\in\mathcal{E}$ with attributes $\mathtt{e}_{u,v}$.
@@
 We use the GATv2 operator Brody et al.
-(2021), which appears to improve static attention problems in graph attention layers.
-Policy gradient methods We consider a standard reinforcement learning setting Sutton and Barto (2018) in which an agent interacts with the environment over a number of steps.
+(2021), which improves static attention problems in graph attention layers.
+Policy gradient methods
+We consider a standard reinforcement learning setting Sutton and Barto (2018) in which an agent interacts with the environment over a number of steps.
 A Markov decision process (MDP) is defined as a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma)$ in which $\mathcal{S}$ is a finite state space, $\mathcal{A}$ is a finite discrete action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ is the transition model, $r:\mathcal{S}\times\mathcal{S}\rightarrow\mathbb{R}$ is the reward function and $\gamma\in(0,1)$ is the discount factor.
@@
 3 Model
-Our proposal for meta-cognitive generalization centers around the ideas of natural decomposition of tasks that an agent can accomplish through interactions with objects.
-Meta-policy recommends localized agent-object interactions.
-The agent executes the interactions following interaction policies and constructs a compositional solution.
-Here, we discuss the main ideas of the framework (see figure 2) and conjecture that it aptly exhibits task generalization characteristics.
-Figure 2: The meta-cognitive generalization framework consists of three levels: meta-level knowledge (K), interaction (I), and execution (X), which are based on levels of abstraction and computation to provide generalization capabilities.
-The knowledge or type space contains type space knowledge graphs representing higher order abstractions.
-The instance space contains object-level instance graphs and functions as an intermediary.
-The execution space is the lowest level abstraction in which the agent acts on meta-level recommendations.
-Computations proceed through recognition, recommendation, execution, and interactions to gather experiences for updates to interaction and meta policies.
-Knowledge An agent can perform low-level actions $u=\{u_{1},u_{2},...\}$ in the environment $\Xi$.
-The environment consists of objects, which the agent’s perception systems associate with the entity types.
-Agent maintains its perceived knowledge in the form of knowledge graphs: a graph of objects and their relationships at the instance level, and a graph of entity types and their abstract relationships at the type level.
-A type space knowledge graph is an ordered pair $G_{K}=(E,\Lambda)$ that consists of a set of entity type nodes, $E=\{E_{1},E_{2},...,E_{m}\}$, and a set of edges or type level relations, $\Lambda=\{\Lambda_{1},\Lambda_{2},...,\Lambda_{\kappa}\}$.
-The same type of relation between objects and abstract entity types has the same meaning; for example, adjacency relations indicate that two objects are adjacent to each other, whether in instance or type space.
-Using the knowledge graph and meta-level action of the meta-policy, $\lambda=\{\lambda_{1},\lambda_{2},...,\lambda_{p}\}$, the meta-level of the agent determines how it will interact with an underlying object.
-To interact with an object, the agent activates an interaction type, $\lambda$, following the meta-policy guidance.
-For an interaction, the agent will perform a series of low-level actions.
-For example, if there is a door, the agent can activate ‘open the door’ at the abstract level, even though it requires many low-level actions such as moving and toggling.
-The agent uses an activation relation to keep track of which objects it wants to interact with and what type of interaction.
-Thus, the agent’s knowledge space $\mathcal{K}=(A,E,\Lambda,\lambda,G_{K})$ comprises of a type space knowledge graph, a set of entity types, abstract relations, and interactions.
-Interaction The environment consists of objects, $e=\{e_{1},e_{2},...,e_{n}\}$, their types, $\text{type}(e_{i})\in E,\text{for}\ e_{i}\in e$, and is partially observable.
-Agent keeps track of the visible objects in its instance space and builds an instance graph, which is an ordered pair of nodes (observed objects) and edges (observed object relations), $G_{I}=(e,\Lambda_{I})$.
-Thus, the agent’s instance space, $\mathcal{I}=(A,e,\Lambda_{I},\lambda,G_{I})$, is made up of its instance space knowledge graph, relationships between objects, and available interactions.
-After constructing the knowledge graphs, the agent’s goal is to identify the objects for interaction which is done by the meta-level.
-Meta-level recommends an interaction goal for the agent.
-For the recommended interaction goals, the agent learns interaction policies for interacting with objects, as well as the relationships between interaction types and entity types.
-With the ability to interact with objects, the agent accumulates interaction knowledge to update the meta-level policy.
-Therefore, when viewing tasks as meta-policy guided interactions with objects, problems are naturally decomposed into interaction goals.
-eXecution By pursuing interaction goals as directed by the meta-level, the agent learns to interact with objects and act at the execution level.
-The execution space for the agent is its partially observed view $X$.
-During an interaction, the agent takes actions $u$ on low-level states $s$.
-After learning both the interaction and meta policies, the agent acts in the environment, aided by interaction with objects.
-Therefore, unlike the models that are tightly coupled with low-level representations through end-to-end policies, agent actions are aided by meta-level knowledge-guided object interactions, providing a mechanism for generalist behaviors.
-Mapping between execution space and instance space Assume that the agent has an object recognition system that can access its partially observed view $X$.
-The agent generates an instance-level graph using the observed objects.
-The function $\Phi:X\to G_{I}$ maps the agent’s object recognition to its instance graph by capturing observed objects and their relationships.
-As the agent interacts with objects, the environment changes, and its recognition system captures a new partially observed view.
-The mapping function then generates the corresponding instance graph.
-Mapping between instance space and type space The function $\Psi:G_{I}\to G_{K}$ provides a mapping between instance and type graphs.
-If the agent encounters a novel object for which it does not have a corresponding typed entity, it must either create a new type in its type space or assign the object to a type that is closely related.
-In the framework, $\Psi$ is learnable and represents a belief of mapping objects to their types.
-However, in this article, we treat the mapping as a fixed mapping from object nodes in $G_{I}$ to their corresponding type nodes in $G_{K}$.
-Proposal for activation object and interaction Using the meta-policy, the agent’s recommendation function $\Upsilon:G_{K},G_{I},\pi\to e_{i},\lambda_{p}$ recommends a candidate object $e_{i}$ for interaction and interaction goal $\lambda_{p}$.
-The recommender iterates through all visible objects in the agent’s instance graph $G_{I}$, then activates the object by adding an activation relationship between the agent and the object to the instance graph, as well as a corresponding activation relation between the agent and the object’s type node in the type space graph $G_{K}$.
-Activated $G_{K}$ represents the activated candidate object and serves as the input state for meta-policy, which returns both the state value and meta-action probabilities.
-The state value represents the value of activating the respective object.
-Thus, the recommender proposes an object for interaction $e_{i}$ by sampling from the values of activating objects, and the proposed interaction $\lambda_{p}$ is sampled from the meta-action probabilities of the proposed object.
-Acting on these recommendations, the agent collects interaction experiences $G_{K},\lambda$, that is, interaction trajectories $G_{K}^{0}\curvearrowright G_{K}^{1}...\curvearrowright G_{K}^{t}$, and learns meta-level policy.
-Interactions At each step, the agent activates the object $e_{i}$ for interaction $\lambda_{p}$, which is its interaction goal, $g^{i}_{p}=(e_{i},\lambda_{p})$.
-Following this interaction goal, the agent interacts with the environment.
-After each round of interaction, the status of the interaction is passed to the meta level.
-Using the meta-policy, the meta-level recommender suggests next target object and interaction.
-The process continues until the task is completed or the environment is terminated.
-Throughout the process, the agent collects experience trajectories at both the interaction and meta levels.
-The policies are then updated based on the collected experiences.
-Meta-policy learning Meta-level representations of type space graphs with activation relations are learned using a message-passing graph neural network.
+Our proposal for metacognitive generalization centers around the ideas of natural decomposition of tasks that an agent can accomplish through interactions with objects.
+The meta-policy recommends localized agent-object interactions, which the agent executes following interaction policies to construct a compositional solution.
+Here, we discuss the main ideas of the framework (see Figure 2) and argue that it exhibits task generalization characteristics.
+Figure 2: The meta-cognitive generalization framework consists of three levels: knowledge ($\mathcal{K}$), interaction ($\mathcal{I}$), and execution ($\mathcal{X}$), which are based on abstraction and computation for providing generalization capabilities.
+The knowledge level contains type space knowledge graphs, which represent higher order abstractions.
+The instance space contains object instance graphs.The interaction level serves as an intermediary, making recommendations for interactions.
+The execution level is the lowest level at which the agent acts based on the recommended interaction subgoals.
+Computations proceed through recognition, recommendation, execution, and interactions to collect experiences for updates to interaction and meta policies.
+Knowledge
+The environment $\Xi$ is partially observable and consists of objects $e=\{e_{1},e_{2},...,e_{n}\}$ and an agent $A$.
+It is assumed that the agent has a recognition system that can access its partial view $X$ and associate objects with their corresponding types, such that the type of an object $e_{i}\in e$ is in entity types $E=\{E_{1},E_{2},...,E_{m}\}$.
+The agent can perform low-level actions $u=\{u_{1},u_{2},...\}$ within the environment.
+When making decisions, the agent organizes its perceived knowledge into knowledge graphs: an instance-level graph and a type-level graph.
+An instance-level graph is an ordered pair $G_{I}=(e,\Lambda)$ of observed objects $e$ and their relations $\Lambda=\{\Lambda_{1},\Lambda_{2},...,\Lambda_{\kappa}\}$.
+A type-level graph is an ordered pair $G_{K}=(E,\Lambda)$ of entity types $E$ and their type level or abstract relations $\Lambda$.
+Relations, whether between objects or entity types, retain the same meaning, such as adjacency indicating that two objects are close to one another.
+The function $\Phi:X\to G_{I}$ maps the agent’s partially observed view to the instance graph by capturing observed objects and deriving their relationships.
+The function $\Psi:G_{I}\to G_{K}$ maps instance space graphs to type space graphs, a direct translation since the agent’s recognition system contains types corresponding to the objects.
+If the agent encounters a novel object without a corresponding type, it must create a new type in its type space or assign the object to a related type.
+In the framework, $\Psi$ is learnable and represents a belief in mapping objects to types.
+However, this article treats the mapping as fixed, from object nodes in the instance graph $G_{I}$ to type nodes in the type graph $G_{K}$.
+As the agent interacts with objects, the environment changes, and its recognition system captures a new partially observed view, resulting in new instance and type space graphs.
+The agent’s type space level is also referred to as the meta-level.
+At the meta-level, the agent uses a meta-policy that accepts a type space knowledge graph and returns meta-level actions or interactions for object interaction.
+Interactions are represented as $\lambda=\{\lambda_{1},\lambda_{2},...,\lambda_{p}\}$.
+Similarly, for each interaction type, the agent uses an interaction policy to guide low-level actions $u$ within the environment.
+Therefore, the agent’s interaction level $\mathcal{I}=(A,G_{I},\lambda)$ comprises of its instance space knowledge graph and available interactions, while the meta-level knowledge $\mathcal{K}=(A,G_{K},\lambda)$ comprises of its type space knowledge graph and interactions.
+Interaction
+After the agent constructs the knowledge graphs, the meta-level recommender $\Upsilon$ recommends a candidate object $e_{i}$ and interaction type $\lambda_{p}$ for an interaction based on the agent’s knowledge graphs and meta-level policy $\pi$.
+The recommender iterates through all visible objects in the agent’s instance graph $G_{I}$.
+For each object, it adds an activation relation between the agent and the object in both the instance space graph $G_{I}$ and the type space graph $G_{K}$, where ‘activating’ refers to observing or focusing on an object for potential interaction.
+For $G_{K}$, the agent’s meta-policy returns the state value and interaction probability.
+The state value represents the value of activating that specific object.
+Then, an object $e_{i}$ for interaction is proposed through sampling based on the values of activation of the visible objects, and interaction $\lambda_{p}$ is sampled from the interaction probability of the proposed object.
+Therefore, at each step, the meta-level of the agent activates a candidate object $e_{i}$ for interaction $\lambda_{p}$, setting the agent’s interaction goal $g^{i}_{p}=(e_{i},\lambda_{p})$.
+Acting on the recommended interaction goal, the agent begins interacting with the environment.
+After each round of interaction, the interaction status is passed to the meta-level, which then proposes the next candidate object and interaction.
+This process continues until the task is completed or the environment is terminated.
+eXecution
+Each agent has a single meta-level policy, whereas interaction policies are a collection of policies corresponding to different interactions.
+By following the interaction goals $g^{i}_{p}$ as directed by the meta-level, the agent interacts with objects.
+During an interaction $\lambda_{p}$, the agent executes a series of low-level actions $u$ on execution level states $s$ based on the relevant interaction policy $\mu_{p}$.
+For example, if there is a door, the agent can activate it and interact with it, like ‘open’ the door, even if this requires multiple low-level actions such as moving and toggling.
+As the agent interacts with objects, it collects trajectories of experiences at both the meta and interaction levels.
+The meta-level trajectory is used to learn the meta-level policy, while the low-level interaction experiences are used to learn the collection of interaction policies.
+The learning approach for the meta-policy and interaction policies is detailed below.
+Meta-policy learning
+Meta-level representations of type space graphs with activation relations are learned using a message-passing graph neural network.
 Nodes have abstract level attributes such as type encoding, while edges have edge type encoding for different types of relationships.
@@
 The model parameters $\theta$ are optimized as $\theta\leftarrow\theta+\alpha_{m}\nabla_{\theta}L_{m}(\theta)$, and the updated meta-policy is used for the next episode of learning.
-Interaction policy learning Interaction policies are a repository of policies for different interactions which can be recommended by the meta-level, $\mu_{\phi}=\{\mu_{\phi_{1}},\mu_{\phi_{2}},...,\mu_{\phi_{p}}\}$ and each policy is parameterized separately.
+Interaction policy learning
+Interaction policies are a repository of policies for different interactions which can be recommended by the meta-level, $\mu_{\phi}=\{\mu_{\phi_{1}},\mu_{\phi_{2}},...,\mu_{\phi_{p}}\}$ and each policy is parameterized separately.
 When an interaction goal with an object $e_{i}$ and interaction $\lambda_{p}$ is recommended, $g^{i}_{p}=(e_{i},\lambda_{p})$, the agent attempts to achieve it by executing a series of low-level actions from $u$.
@@
 For each interaction policy, the model parameters $\phi_{p}$ are optimized as $\phi_{p}\leftarrow\phi_{p}+\alpha\nabla_{\phi_{p}}L_{p}(\phi_{p})$, and the updated interaction policies are used for the next episode of learning.
-As a result, the agent follows a trajectory of activated interactions with objects and aids the formation of a task solution through meta-guided interactions.
-The proposed framework is generic to be applied to a wide range of domains.
-4 Experiments
-Experimental set up To assess the capabilities of the proposed agents, we require a challenging environment with ingredients for higher order reasoning to solve tasks.
-Our goal here is to investigate whether the introduction of meta-knowledge guided interactions benefits agents.
+In summary, when viewing tasks as meta-policy guided interactions with objects, problems are naturally decomposed into interaction goals.
+Unlike the models that are tightly coupled with low-level representations through end-to-end policies, in the proposed framework, agent actions are aided by meta-level knowledge guided object interactions, providing a mechanism for generalist behaviors.
+4 Results and Discussion
+Experiments
+Our goal is to examine whether meta-level knowledge-guided interactions benefit agents in task generalization.
+To assess the capabilities of the proposed agents, we need a challenging environment having ingredients for higher-order reasoning to solve tasks.
 MiniGrid world environments Chevalier-Boisvert et al.
-(2018) are suitable for our purpose.
-These are procedurally generated environments in which the agent interacts with objects such as doors, keys, balls, and boxes.
-Figure 3: MiniGrid ObstructedMaze Full environment is an appropriate testbed because it is partially observable, requires complex reasoning, and new tasks can be created.
-The environments are challenging due to partially observability, sparsity of rewards and random goal locations.
-The MiniGrid-ObstructedMaze-Full environment, figure 3, in particular, is a difficult environment with only a few claims for solving it.
-Due to the large state space size, stochasticity and partial observability, agents cannot memorize but must learn.
-In the environment, the agent is initially placed in the middle room, with a 7x7 view.
+(2018) are suitable for this purpose; they are procedurally generated environments where the agent interacts with objects such as doors, keys, balls, and boxes.
+ObstructedMaze-Full
+ObstructedMaze-1Dlhb
+Figure 3: Environments: ObstructedMaze-Full is used to experimentally evaluate the agent’s performance on the designed tasks, while ObstructedMaze-1Dlhb is used to evaluate the transferability of interaction policies.
+The obstructed maze environments, shown in Figure 3, are particularly challenging due to partial observability, stochasticity, and sparse rewards, with only a few claims for solving it.
+In the MiniGrid-ObstructedMaze-Full environment, the agent is initially placed in the middle room, with a 7x7 view.
 At the low level, the agent can move forward, left, right, pick, drop, and toggle, and its goal is to pick up the blue ball, which is randomly placed in one of the corner rooms.
@@
 Keys are hidden inside boxes, but the agent cannot see inside the boxes without toggling.
-The original task is referred as Task 0.
-Furthermore, we identify the following additional task scenarios that were not seen during training but were presented to the agent during testing:
+The default task is referred to as Task 0.
+Furthermore, we identify the following new task scenarios, which were presented to the agent during testing but not during training:
   • No context in training (Task 1) - a relation that the agent did not come across during training.
@@
 To simulate, as soon as the agent enters the room with the goal, relocate the goal to one of the neighboring rooms.
-Agent training
-Initialize model weights $\theta,\phi$ , parameters, $T$ for  $t=1$ to $T$ do
-       // collect exp for meta exp size K
+It is only done once, not repeatedly.
+Agents learning to interact using type space knowledge
+initialize model weights $\theta,\phi$ , parameters, $T$
+for  $t=1$ to $T$ do
+       /* collect exp for meta exp size K */
        for $k=1$ to $K$ do
-             make instance graph $G_{I}\leftarrow\Phi(X)$ map $G_{I}$ to $G_{K}$, $G_{K}\leftarrow\Psi(G_{I})$ proposal $(e_{i},\lambda_{p})\leftarrow\Upsilon(G_{K},G_{I},\pi)$ // 1st variant
-             interaction exp following $\mu_{\lambda_{p}}$ // 2nd variant
-             reaching exp following $\mu_{\text{reach}}$ if success then
-                   interaction exp following $\mu_{\lambda_{p}}$
-             end if
-            collect exp $\mathcal{D}_{\tau_{p}}$ and compute $A_{p}$
+             observe a partial view $X$
+             make instance graph $G_{I}\leftarrow\Phi(X)$
+             map $G_{I}$ to $G_{K}$, $G_{K}\leftarrow\Psi(G_{I})$
+             recommend $(e_{i},\lambda_{p})\leftarrow\Upsilon(G_{K},G_{I},\pi)$
+             interact with object $e_{i}$ following $\mu_{\lambda_{p}}$
+             collect interaction exps $\mathcal{D}_{\tau_{p}}$ and compute $A_{p}$
        end for
-      collect meta exp $\mathcal{D}_{\tau_{m}}$ and compute $A_{m}$ // Update policies, both variants
-       update $\pi$ by optimizing $\mathcal{L}_{m}(\theta)$ for  $\lambda_{p}$ do
+      collect meta exp $\mathcal{D}_{\tau_{m}}$ and compute $A_{m}$
+       /* Update policies */
+       update $\pi$ by optimizing $\mathcal{L}_{m}(\theta)$
+       for  $\lambda_{p}$ do
              update $\mu_{\lambda_{p}}$ by optimizing $\mathcal{L}_{p}(\phi_{p})$
        end for
-      // 2nd variant only
-       update $\mu_{\text{reach}}$ by optimizing $\mathcal{L}_{\text{reach}}(\phi_{\text{reach}})$
 end for
-Algorithm 1 KIX
-We devise two variants, KIX1 and KIX2.
-Both variants learn meta-policy and interaction policies.
-However the second variant additionally learns a reachability policy, so it interacts with an object once it has reached it.
-All meta, interactions, and reachability policies are implemented as actor-critic PPO policies.
-Algorithm 1 provides a high-level overview of the implementation.
-The input state embeddings for interaction and reachability actor-critic networks are constructed by a convolutional network with three Conv2D layers with 16, 32, and 64 output channels, a max pooling layer added after the first convolution layer, and exponential linear units used as activation functions.
-The meta-level recommends various objects for interaction, so an activation indicator for the object of interest is added to a new input channel, creating a four-channel input.
-Actors and critics are fully connected networks, each with one hidden layer.
-Meta-level policies are learned over type space knowledge graphs.
-In a type space graph, node object type information is encoded as an attribute vector.
-Edges in type space graph represent type-level relationships between nodes, such as "visible," "adjacent," "carrying," and "activated." A one-hot encoded edge attribute vector is used to encode these relationships.
-Embeddings for meta-level actor-critic are provided by two graph attention convolution layers, each with 16-dimensional embeddings and four attention heads.
+Algorithm 1 KIX procedure
+We devised two variants: KIX-A and KIX-R.
+Both variants learn meta-policy and a collection of interaction policies.
+While the KIX-A variant combines both reaching objects and interacting with them, the KIX-R variant learns an additional interaction policy referred to as the ‘reachability’ policy, which is used to reach objects prior to interacting with them.
+All policies, including meta and interactions, are implemented as actor-critic PPO policies.
+Algorithm 1 provides an overview of the procedure.
+The agent’s partially observed view $X$ contains objects with identifiers and attributes like color and type.
+The function $\Phi$ takes $X$ and creates an instance space graph by using visible objects as nodes, including the agent, and determining relationships between nodes such as ‘visible’, ‘adjacent’, and ‘carrying’.
+The ‘adjacent’ relation is derived using distance between the two objects.
+When the agent chooses an object to interact with, an ‘activated’ relation is added between the agent and the object node.
+The function $\Psi$ then constructs a type space graph from the instance space graph, encoding node types as attribute vectors.
+Edges in a type space graph have the same meaning as those in an instance space graph and are encoded as one-hot encoded edge attribute vectors.
+Figure 4: Agents learning to interact: The learning progress of the agents KIX-A (blue) and KIX-R (red) is shown as returns over steps, which illustrate both the meta-level and interactions such as ‘open’, ‘open with key’, ‘reveal’, ‘pickup’ (not shown for KIX-R), ‘drop’, and ‘reach’ in the ObstructedMaze-Full environment.
+Agents learn how to perform low-level actions in the execution space using interaction policies.
+A convolutional network with three Conv2D layers with 16, 32, and 64 output channels, a max pooling layer added after the first convolution layer, and exponential linear units as activation functions is used to provide the input state embeddings for the interaction actor-critic.
+An activation indicator for the proposed interaction object is added as a new input channel, resulting in a four-channel input to the filters.
+Actor-critic networks are fully connected with a single hidden layer.
+Agents learn a meta-policy in the type space.
+Two graph attention layers with 4 heads and 16 dimensional embedding provide 64 dimensional input embeddings for the meta policy.
 During message passing, these layers share weights between source and target.
-The final graph attention convolution layer is used for graph readout with a global add pool.
-Meta-level actors and critics are fully connected networks, with the critic learning a single value and the actor learning the probability of taking a meta-level action, such as ’pickup’, ’drop’, ’reveal’, ’open’, or ’open with key’.
-The meta-level actions serve as interaction goals.
-The environment rewards the agent with $r=1-(c/c_{T})$ where $c$ is the number of steps taken and $c_{T}$ is the max possible steps.
-For interaction policy learning, the agent receives a reward of 1 when it has a successful interaction.
-At the meta-level it is rewarded with 0.1 for a successful proposed interaction and $1+r$ for achieving the environment goal.
-A base agent is also implemented, with a similar network and a PPO policy, and it receives environmental rewards.
-Grid search is used to select hyperparameters.
-All the agents are trained for the same number of environment steps using 32 parallel workers.
-Figure 4: The return profiles of agents performing various evaluation tasks show that the proposed agents have higher returns on all tasks and benefit from meta-guided interactions.
-Meta-guided interactions aid agents Our proposal is motivated by the idea that, rather than acting directly at the execution level, meta-level guided interactions with objects aid agents in generalizing while performing tasks.
-Because the agent’s environmental return is determined by the number of steps taken to complete the task, completing tasks in fewer steps yields a higher return.
-Therefore, the distribution of environmental returns is a reliable indicator of whether meta-level interaction recommendations benefit agents.
-We evaluate the distribution of environmental returns for each agent across all four tasks.
-For each task, the trained agents are run for 12K episodes by parallel workers, and their environmental rewards are recorded.
-In order to understand the effects of meta-guided interactions and to ensure a fair comparison between agents where differences in success rates are not a factor, we examine the top $k$ returns with $k=100$.
-Figure 4 illustrates the return profiles of agents.
-It is clear from the distribution of returns that the proposed agents have received a higher return on all tasks than the base agent, furthermore, this is even more so for dynamic tasks like task 3.
-It demonstrates that meta-guided interactions not only help agents complete their task goals in fewer steps, but also aid them to perform a wider range of tasks in static or dynamic environments.
-Furthermore, the higher returns of the second variant demonstrate that agents benefit from the introduction of reachability as a separate interaction.
-In the absence of a separate reachability policy, reaching is implicit in each interaction policy; however, separating allows for a shared learning of reachability followed by interactions which could be beneficial particularly in dynamically changing environments.
-Figure 5: Wasserstein distances between traversal distributions, when agents complete tasks 1-3, and the trained task 0, provide comparative insights.
-Agents profile While performing the tasks, agents interact with objects in the environment according to the recommended interaction policies; as a result, agents navigate and visit rooms.
-Because interactions are recommended by the meta-level, which is only trained on the training task, tracking agents’ room visits across all tasks provides an insight about their learning.
-Let $p_{0}$ be the distribution of agent room visits for trained task 0, and $p_{i},i\in\{1,2,3\}$ be the distributions for tasks 1–3.
-For both variants of the agent, we then compute the Wasserstein distance, a metric from optimal transport theory, between the visit distribution for tasks 1-3 and the trained task 0, as shown in figure 5.
-The distribution of visits to each agent is expected to be different from the training task, due to changes in the evaluation tasks.
-However, our interest is in how differently the two variants behave in evaluation tasks, as indicated by the computed metric.
-In task 1, the goal room remained unchanged, but a new relation was added to the evaluation.
-For both variants, the computed metric value for task 1 is similar, indicating that they transfer their learning to this task in a similar manner.
-However, there is a noticeable difference for tasks 2 and 3.
-In task 2, doors in the central room are locked with keys kept in the room - a static task, whereas in task 3, the goal is dynamically relocated based on observance - a dynamic task.
-The computed metric for the dynamic task for the second variant is smaller than the first variant, implying that its evaluation visit distribution is closer to that of the training task.
-It suggests that the learning in the second variant can better capture dynamic changes.
-When combined with the return profiles, it reinforces that having an additional reachability policy benefits the agent in a wider range of tasks, particularly dynamic tasks, and improves generalization.
+The final graph attention layer is then used for graph readout, applying a global add pool.
+The meta-level actor and critic are both fully connected networks, with the critic learning a single value and the actor learning the likelihood of taking the meta-level actions ‘pickup’, ‘drop’, ‘reveal’, ‘open’, or ‘open with key’, each of which corresponds to an interaction policy.
+Table 1: Training hyper-parameters
+Parameter Value Parameter Value
+meta exp size 8 gae lambda 0.95
+exp size 64 entropy coeff 1e-04
+meta minibatch size 64 value coeff 1.0
+minibatch size 64 meta entropy coeff 1e-04
+meta learning rate 1e-05 meta value coeff 1.0
+learning rate 1e-05 ppo clip 0.2
+gamma 0.99 ppo steps 4
+The environment rewards the agents with $r=1-(c/c_{T})$ where $c$ is the number of steps taken and $c_{T}$ is the max possible steps.
+For interaction policies, the agent receives a reward of 1 when it has a successful interaction.
+At the meta-level, it is rewarded with 0.1 for a successful proposed interaction and $1+r$ for achieving the environment goal.
+The base agent, BASE, has a similar actor-critic architecture and receives environment rewards.
+The training hyper-parameters are given in Table 1.
+Learning rates were determined using grid search with values 1e-3, 1e-4, 1e-5, and 1e-6.
+All the agents are trained for $\sim 6.2\times 10^{8}$ steps, using 32 parallel workers.
+As shown in Figure 4, as training progresses, agents learn to interact with objects and associate higher-level structured knowledge with these interactions using object types and knowledge graphs, as indicated by higher meta-level returns.
+This, in turn, leads to better interaction experiences and further improvements in interaction policies.
+Furthermore, the KIX-R variant first focuses on learning how to reach different objects, but once it learns to reach, it begins to learn interactions in fewer steps and with much higher returns compared to the KIX-A variant.
+It suggests that the agents learn higher-level relations and apply these learned relations and interactions with objects to solve problems.
+Meta-level knowledge-guided interactions aid agents
+Our approach is motivated by the idea that meta-level guided interactions with objects help agents in generalizing while performing tasks, rather than acting directly at the execution level.
+Since the agent’s environmental return is determined by the number of steps taken to complete a task, completing tasks in fewer steps yields a higher return.
+Therefore, the distribution of environmental returns is a reliable indicator of the value of meta-level interaction recommendations for agents.
+We evaluate the distribution of environmental returns for each agent across four tasks.
+For each evaluation task, trained agents are run for 12K episodes by parallel workers, and their environmental rewards are tracked.
+Figure 5: Evaluation return profiles of agents performing various tasks show that the proposed agents benefit from meta-level guided interactions.
+To understand the effects of meta-guided interactions and ensure a fair comparison between agents, regardless of success rate differences, we examine the top $k$ returns with $k=100$.
+Figure 5 shows the return profiles of agents.
+It is evident from the distribution of returns that the proposed agents achieve higher returns on all tasks compared to the base agent, including dynamic tasks like Task 3.
+This demonstrates that meta-guided interactions not only assist agents in completing tasks in fewer steps, but also enable them to perform a wider range of tasks in both static and dynamic scenarios.
+Furthermore, higher returns of the second variant show that agents benefit from treating reachability as a separate interaction.
+In the absence of separate reachability, reaching is embedded into interactions.
+Separate reachability allows shared learning followed by interactions, which is particularly beneficial in dynamically shifting scenarios.
+Transferability of policies
+(a) Meta training
+(b) Evaluation task returns
+Figure 6: Evaluation return profiles show that interaction policies can be transferred from one environment to another.
+The ObstructedMaze-1Dlhb environment is used to assess the transferability of interaction policies learned in one environment to another.
+In this environment, the agent’s goal is to pick up a blue ball hidden in a two-room maze.
+The rooms are separated by a door, which is obstructed by a ball and locked.
+The keys are hidden inside boxes.
+Agents employ previously acquired interaction policies from the base task in the ObstructedMaze-Full environment while only learning the meta-policies.
+During evaluation, agents use the learned meta-policy, but the interaction policies come from the ObstructedMaze-Full environment itself.
+Figure 6 shows the meta training returns and evaluation return profiles, similar to the main environment, for both agent variants.
+It is evident that agents have gained the ability to exploit interaction policies from the previous environment, and the second variant, which employs a separate reachability policy, appears to have benefited more.
 5 Conclusion
 In this work, we proposed a metacognitive framework that promotes generalist behaviors through three main contributions.
-First, generalist and flexible behaviors necessitate abstraction and composition, but current models lack natural abstractions and compositions that can preserve both object abstractions and relational knowledge structures.
-In our model, natural abstractions and compositions are represented and grounded using type space knowledge graphs which also preserve relational knowledge between objects.
-Second, rather than directly acting in the raw environment, interactions with objects play an important role in how an agent navigates the environment to complete tasks.
-To accomplish this, we propose an intermediary interaction level.
-The meta-level of the agent learns from type space knowledge graphs and recommends what objects to interact with, as well as what interaction should be.
-In this sense, our model also functions as an interaction goals generating machine.
-Third, interaction policies learned through interaction goals serve as a repository of policies for the agent to use in the execution space, providing for the necessary loose coupling for generalization via interactions as intermediaries.
-The agent’s meta-level recommends interactions based on knowledge structures and learned meta-policy, and the execution level applies interaction policies to complete tasks.
-We demonstrated that metacognitive knowledge guided interactions aid agents in generalization.
-Our framework is modular and adaptable, with the potential to have implications across domains through the induction or integration of knowledge structures, whether extrinsic, intrinsic, or derived, as well as control variables at the meta or interaction level.
+First, generalist behaviors necessitate abstraction and composition; however, current models lack natural abstractions and compositions that can preserve both object abstractions and relational knowledge structures.
+In our model, type space knowledge graphs are used to represent and ground natural abstractions and compositions while also preserving relational knowledge between objects.
+Second, rather than directly acting in the environment, interactions with objects are critical in performing the tasks.
+To accomplish this, introduce an intermediary interaction level.
+The agent’s meta-level learns from type space knowledge graphs and makes recommendations on which objects to interact with and how to interact with them.
+In this sense, our model functions as an interaction goal-generating machine.
+Third, interactions employing interaction policies enable the necessary loose coupling required for generalization by acting as intermediaries.
+We demonstrated that metacognitive knowledge-guided interactions aid agents in generalization.
+Lastly, Our framework is modular and adaptable, allowing for the integration of knowledge structures and control variables at both the meta and interaction levels.
+This framework is promising, particularly for domains where higher-order knowledge and interaction compositions can be exploited, such as robotics, autonomous and cognitive systems.

### v3

@@
 Abstract
-People aptly exhibit general intelligence behaviors in solving a variety of tasks with flexibility and ability to adapt to novel situations by reusing and applying high-level knowledge acquired over time.
-But artificial agents are more like specialists, lacking such generalist behaviors.
-Artificial agents will require understanding and exploiting critical structured knowledge representations.
-We present a metacognitive generalization framework, Knowledge-Interaction-eXecution (KIX), and argue that interactions with objects leveraging type space facilitate the learning of transferable interaction concepts and generalization.
-It is a natural way of integrating knowledge into reinforcement learning and is promising to act as an enabler for autonomous and generalist behaviors in artificial intelligence systems.
-1 Introduction
-Figure 1: Learning generalist behaviors requires developing a natural and abstract understanding of the environment, as well as meta-reasoning and interaction abilities.
-The metacognitive knowledge-guided approach exploits inherent abstractions and relationships between objects to generate interaction goals, enabling the agent to interact with objects and generalize to solve complex tasks.
-Flexibility, ability to actively think and evaluate options –learning and reasoning to enable adaptation– are hallmarks of human intelligence.
-Core to this intelligence is how easily people can find natural decomposition in world environments and tasks; this enables flexible, generalist behaviors.
-Replicating such autonomy and general intelligence abilities in artificial agents continues to pose significant challenges, which we argue requires a knowledge and interaction-centric metacognitive approach.
-Meta-reasoning about next actions, their values, and possible next states facilitates flexible decision-making in novel situations, problem solving, concept learning, and knowledge acquisition.
-This reasoning extends beyond mere memory lookup.
-Meta-reasoning is more broadly sense-making in a knowledge system - enabling people to intuitively direct their learning to make sense of relationships between entities, interact with the same entities in diverse ways, and ground their interactions for both short- and long-term goals.
-Metacognitive reasoning appears to be central to generalization.
+People aptly exhibit general intelligence behaviors through flexible problem-solving and the ability to adapt to novel situations by reusing and applying high-level knowledge acquired over time.
+In contrast, artificial agents tend to be specialists, lacking such generalist behaviors.
+To bridge this gap, artificial agents will require understanding and exploiting critical structured knowledge representations.
+We introduce a metacognitive reasoning framework, Knowledge-Interaction-eXecution (KIX), and argue that interactions with objects, by leveraging a type space, facilitate the learning of transferable interaction concepts and promote generalization.
+This framework offers a principled approach for integrating knowledge into reinforcement learning and holds promise as an enabler for generalist behaviors in artificial intelligence, robotics, and autonomous systems.
+Introduction
+The ability to actively reason, flexibly evaluate options and learn from experience is a hallmark of human intelligence.
+Core to this intelligence is the human ability to identify natural decompositions within complex environments and tasks with ease, enabling flexible and generalist behaviors.
+Replicating such autonomy and general intelligence abilities in artificial agents continues to pose significant challenges, which we argue necessitate a knowledge and interaction-centric metacognitive approach.
+Meta-reasoning about next actions, their values, and possible next states underpins flexible decision-making in novel situations, problem solving, concept learning, and knowledge acquisition.
+This form of reasoning extends beyond mere memory lookup.
+More broadly, meta-reasoning functions as a sense-making process within a knowledge system, enabling people to intuitively direct their learning to discern relations between entities, interact with the same entities in diverse ways, and ground these interactions for both short- and long-term goals.
+Metacognitive reasoning thus emerges as a foundational mechanism for generalization.
 In contrast, meta-reasoning systems in artificial intelligence are primarily focused on resource allocation problems.
-While resource allocation underpins efficient functioning, it is crucial to ideate metacognitive generalization, Figure 1, to enable generalist behaviors in artificial agents.
-Our proposal builds on three main principles for solving complex tasks flexibly.
-Natural Decomposition and Abstraction: World environments are naturally composed of entities categorized by their types, similar to a natural hierarchy of objects and classes.
-Agents interact with entities in the instance space, which inherit properties and relationships from a higher-level type space.
-Interactions: World environments are essentially knowledge graphs representing underlying objects and their relationships.
-An agent’s metacognitive process employs these knowledge graphs to create a plan specifying which entities to interact with and the nature of these interactions to complete tasks and achieve goals.
-Composition of Interactions: Agents’ flexible behaviors stem from their ability to learn interactions and leverage interaction pathways for graph traversal at both the type level and the instance level, enabling quick adaptation.
+While resource management is essential for efficient functioning, it is crucial to ideate metacognitive generalization, illustrated in Figure 1, to enable generalist behaviors in artificial agents.
+(a) Overview
+(b) Approach
+Figure 1: a) Illustration of metacognitive reasoning process in a navigation domain, where an agent (depicted as a red triangle) must interpret its environment to achieve the goal of retrieving a blue ball located behind a locked door and various obstacles.
+This process requires reasoning over abstract object types and their relations, which in turn guide the agent’s interactions with objects.
+Such abstract reasoning fundamentally relies on higher-order abstract structures and compositional interactions over these abstractions.
+b) Learning generalist behaviors requires developing a natural and abstract understanding of the environment, as well as meta-reasoning and interaction abilities.
+The metacognitive knowledge-guided approach exploits inherent abstractions and relations between objects to generate interaction goals, enabling the agent to interact with objects and generalize to solve complex tasks.
+Metacognitive reasoning is fundamentally grounded in interaction, requiring agents to draw upon their interpretations of abstract knowledge to guide their interactions.
+Interpretation precedes and shapes interaction; it involves developing both natural and abstract understandings of the environment that can be deployed as needed to make sense of the current state and plan subsequent interactions.
+To illustrate this process, consider the example in Figure 1a, which depicts an agent and its interactions in a navigation domain.
+To build knowledge of the environment, an agent must develop a conceptual understanding of the objects in the environment and how to interact with them.
+As some objects may even appear similar, the agent must decipher their abstract types.
+Learning to interact then involves learning which interactions are possible with these abstract object types and the relations between them.
+Type serves as a natural abstraction, and interpretation thus incorporates this meta-level reasoning over abstract knowledge.
+The utility of interactions is then learned through meta-level policies in this abstract space, which can subsequently be leveraged to recommend interactions and execute compositions of interactions at a lower level, thereby facilitating problem-solving.
+Our proposal is based on three main principles for flexibly solving complex tasks.
+Natural abstraction: Real-world environments are naturally composed of entities categorized by their types, similar to a natural hierarchy of objects and classes.
+Entities in the instance space inherit properties and relations from a higher-level type space.
+Type-space graphs represent the underlying objects and their relations in environments.
+This abstraction is where the knowledge resides and it is the knowledge or meta level.
+Interactions: An agent’s metacognitive process employs the type-space graphs to create a plan specifying which entities to interact with and the nature of these interactions to complete tasks and achieve goals.
+While the metacognitive policy guides agents based on abstract knowledge, agents interact with entities in the instance space which are instantiations of abstract types.
+This is the interaction level.
+Composition of interactions: The agent interacts with meta-guided objects as a recommended interaction type and executes low-level policies corresponding to a guided interaction type.
+Therefore, a composition of these interactions as interaction pathways for graph traversal at both the meta level and the instance level facilitates execution of low-level actions which are loosely coupled with abstract knowledge through interactions.
 Integrating these principles into metacognitive learning, reasoning, and control provides the levers and connecting points for instilling generalist behaviors in agents.
+Agents’ flexible behaviors stem from their ability to interpret knowledge, learn how to interact with objects through meta level policies and leverage interaction pathways enabling quick adaptation.
 The proposed metacognitive framework captures the essence of generalist behaviors and may have far-reaching implications across various domains.
-In this article, we describe the key components of the framework that exemplify generalist characteristics, followed by an experimental evaluation as well as how we envision it enabling flexibility through meta-level knowledge guided interactions.
-2 Background
+We describe the key components of the framework that exemplify generalist characteristics, followed by an experimental evaluation in a navigation domain and discuss how we envision it enabling flexibility through meta-level guided interactions.
+Background
 In this section, we briefly describe closely related work and the general setting of meta-reasoning, graph neural networks, and policy gradient methods.
@@
 Compositionality and abstractions are central to generalization learning, with abstraction requiring a mechanism for grounding representations in actual objects.
-In previous work, Veerapaneni et al.
-(2020) presented an object-centric model where object representations were constructed from raw observations and simpler local functions with uniformity, which were then used for prediction and planning.
-Similarly, Vijay et al.
-(2019) used a symbolic object-centric representation to study the impact of adding prior knowledge through relations in an object graph.
+An object-centric model Veerapaneni et al.
+(2020) was presented in which object representations were constructed from raw observations and simpler local functions with uniformity, which were then used for prediction and planning.
+Similarly, a symbolic object-centric representation Vijay et al.
+(2019) was used to study the impact of adding prior knowledge through relations in an object graph.
 These studies focus primarily on object space, which inhibits the abstraction of objects and relational structures naturally occurring in the real world and used for decision-making.
@@
 We address these largely overlooked questions by preserving abstract knowledge and relational higher-order structures in the form of knowledge graphs, where abstractions are grounded using type space.
-Learning a compositional structure is achieved through learning over knowledge graphs, with neighborhood aggregation until stability is reached.
-This distinguishes our approach from previous works.
-Other methods emphasize different aspects, such as encouraging exploration through intrinsic reward or information gain Ten et al.
+Learning a compositional structure is achieved through learning over knowledge graphs with neighborhood aggregation.
+Hierarchical reinforcement learning approaches Dayan and Hinton (1992); Sutton et al.
+(1999); Pateria et al.
+(2021); Bagaria et al.
+(2021) learn a higher-level policy by choosing subtasks as higher-level actions, which decompose tasks into subtasks.
+Higher-level actions persist over a longer timescale, referred to as temporal abstraction Barto and Mahadevan (2003); Dietterich (2000); Sutton et al.
+(1999).
+In contrast to options Sutton et al.
+(1999), which combine a subtask with the core MDP, MAXQ value function decomposition Dietterich (2000) decomposes the core MDP into smaller sub-MDPs, each with a separate learnable policy.
+Other methods induce learning hierarchies based on teacher signals Portelas et al.
+(2020); Hutsebaut-Buysse et al.
+(2022) or a progression from easier to more difficult tasks Campero et al.
+(2020).
+These methods enable agents to learn hierarchy of policies, but they remain in the object space for both low- and high-level actions and do not meet a key meta-reasoning demand: type abstractions and loose coupling between abstraction and execution space.
+Methods such as Ten et al.
 (2022); Oudeyer et al.
-(2007); Pathak et al.
-(2017), motivation Moerland et al.
-(2023); Zhu et al.
-(2023), developing internal goals for empowerment Klyubin et al.
-(2005), using a curriculum to progress from easier to more difficult tasks Soviany et al.
-(2022); Campero et al.
-(2020), or inducing a learning hierarchy based on teacher signals Portelas et al.
-(2020) Pateria et al.
-(2021); Hutsebaut-Buysse et al.
-(2022) or rewards Furelos-Blanco et al.
-(2023).
-These methods focus on the execution space, influencing reward signals for learning policies, but largely ignore the higher-order structures and relations between objects in the environment.
-Our work distinguishes itself by incorporating natural relational structures through knowledge graphs and leverages meta-level knowledge structures to guide interactions that govern low-level execution.
+(2007) encourage exploration through intrinsic reward or motivation, Klyubin et al.
+(2005) develop internal goals for empowerment, focusing on aspects different than ours.
+On the other hand, neuro-symbolic reasoning approaches Hitzler and Sarker (2022); Zhang and Sheng (2024) aim to overcome the limitations of purely symbolic or purely neural methods by integrating them into hybrid models that combine the learning capacity of neural networks with the logical inference and interpretability of symbolic systems.
+However, significant challenges persist in facilitating interaction and integration between these disparate paradigms to ensure a seamless, flexible, and semantically grounded natural unification.
+Our approach meets these demands by leveraging natural higher-order knowledge structures in the form of type space graphs for meta-level policy and facilitating interactions with objects via an intermediary between meta-level and execution space, thus enabling loose coupling and reuse of low-level policies.
+The appeal of our approach is that interaction policies can be swapped in and out depending on the interaction type.
+Our work distinguishes itself by incorporating relational abstract knowledge structures to guide interactions that govern low-level execution.
 Meta-reasoning
@@
 The monitoring process allows the meta controller to track the status of object-level processes and actions, as well as the consequences of those actions.
-However, current meta-cognitive reasoning, primarily as a resource manager, is insufficient for exhibiting generalist behaviors, necessitating generalist meta-cognition approaches such as the one proposed here.
+However, current meta-reasoning, primarily as a resource manager, is insufficient for exhibiting generalist behaviors, necessitating generalist meta-cognitive approaches such as the one proposed here.
 Graph neural networks
@@
 Typically, the representation is learned through message passing or neighborhood feature aggregation, in which node embedding vectors are updated in each layer based on message aggregation from neighbors:
-$h_{v}^{(k)}=\mathcal{F}^{\prime}\left(h_{v}^{(k-1)},\bigoplus_{u\in\mathcal{N}(v)}\mathcal{F}^{\prime\prime}\left(h_{v}^{(k-1)},h_{u}^{(k-1)},\mathtt{e}_{u,v}\right)\right)$
-where $h_{v}^{(k)}$ is the updated representation of node $v$ embedding in layer $k$, $\mathcal{N}(v)$ is the set of neighboring nodes of node $v$, $h_{u}^{(k-1)}$ is the representation of neighboring node $u$ in layer $k-1$, $\bigoplus$ is a differentiable, permutation invariant aggregation function, $\mathcal{F}^{\prime}$ and $\mathcal{F}^{\prime\prime}$ are differentiable functions usually neural networks.
+$h_{v}^{(k)}=\mathcal{F}^{\prime}\left(h_{v}^{(k-1)},\bigoplus_{u\in\mathcal{N}(v)}\mathcal{F}^{\prime\prime}\left(h_{v}^{(k-1)},h_{u}^{(k-1)},\mathtt{e}_{u,v}\right)\right)$ (1)
+where $h_{v}^{(k)}$ is the updated representation of node $v$ embedding in layer $k$, $\mathcal{N}(v)$ is the set of neighboring nodes of node $v$, $h_{u}^{(k-1)}$ is the representation of neighboring node $u$ in layer $k-1$, $\bigoplus$ is a differentiable permutation invariant aggregation function, $\mathcal{F}^{\prime}$ and $\mathcal{F}^{\prime\prime}$ are differentiable functions usually neural networks.
 Node features are transformed prior to aggregation, and the methods used to transform and aggregate features differ between graph network variants.
@@
 At each step t, the agent takes an action $a_{t}\in\mathcal{A}$ in state $s_{t}\in\mathcal{S}$ according to its policy $\pi$, which is a mapping of states to actions.
-As a result, the agent reaches state $s_{t+1}\in\mathcal{S}$, receives a scalar reward $r_{t}\in\mathbb{R}$, and accumulates returns; $R_{t}=\sum_{k=0}^{\inf}\gamma^{k}r_{t+k+1}$ represents the accumulated return from step t.
+As a result, the agent reaches the state $s_{t+1}\in\mathcal{S}$, receives a scalar reward $r_{t}\in\mathbb{R}$, and accumulates returns; $R_{t}=\sum_{k=0}^{\inf}\gamma^{k}r_{t+k+1}$ represents the accumulated return from step t.
 The agent’s goal is to maximize the expected return for each state.
@@
 Proximal policy optimization (PPO) Schulman et al.
-(2017) is an on-policy policy optimization technique that addresses concerns in actor critic methods using clipped objectives.
+(2017) is an on-policy policy optimization technique that addresses concerns in actor-critic methods using clipped objectives.
 We adapt PPO for learning policies.
-3 Model
-Our proposal for metacognitive generalization centers around the ideas of natural decomposition of tasks that an agent can accomplish through interactions with objects.
+Metacognitive framework
+Our proposal for metacognitive generalization centers around the ideas of natural decomposition of tasks that an agent can accomplish through interpretation of knowledge and interpretation guided interactions with objects.
 The meta-policy recommends localized agent-object interactions, which the agent executes following interaction policies to construct a compositional solution.
-Here, we discuss the main ideas of the framework (see Figure 2) and argue that it exhibits task generalization characteristics.
+Here, we discuss the main ideas of the framework (see Figure 2) and argue that it exhibits task-generalization characteristics.
 Figure 2: The meta-cognitive generalization framework consists of three levels: knowledge ($\mathcal{K}$), interaction ($\mathcal{I}$), and execution ($\mathcal{X}$), which are based on abstraction and computation for providing generalization capabilities.
 The knowledge level contains type space knowledge graphs, which represent higher order abstractions.
-The instance space contains object instance graphs.The interaction level serves as an intermediary, making recommendations for interactions.
+The instance space contains object instance graphs.
+The interaction level serves as an intermediary, making recommendations for interactions.
 The execution level is the lowest level at which the agent acts based on the recommended interaction subgoals.
@@
 The agent can perform low-level actions $u=\{u_{1},u_{2},...\}$ within the environment.
-When making decisions, the agent organizes its perceived knowledge into knowledge graphs: an instance-level graph and a type-level graph.
+When making decisions, the agent organizes its perceived knowledge into graphs: an instance-level graph and a type-level graph.
 An instance-level graph is an ordered pair $G_{I}=(e,\Lambda)$ of observed objects $e$ and their relations $\Lambda=\{\Lambda_{1},\Lambda_{2},...,\Lambda_{\kappa}\}$.
@@
 As the agent interacts with objects, the environment changes, and its recognition system captures a new partially observed view, resulting in new instance and type space graphs.
-The agent’s type space level is also referred to as the meta-level.
-At the meta-level, the agent uses a meta-policy that accepts a type space knowledge graph and returns meta-level actions or interactions for object interaction.
+The agent’s knowledge level is also referred to as the meta-level.
+At the meta-level, the agent uses a meta-policy that accepts a type space graph and returns interactions or meta-level actions for object interaction.
 Interactions are represented as $\lambda=\{\lambda_{1},\lambda_{2},...,\lambda_{p}\}$.
 Similarly, for each interaction type, the agent uses an interaction policy to guide low-level actions $u$ within the environment.
-Therefore, the agent’s interaction level $\mathcal{I}=(A,G_{I},\lambda)$ comprises of its instance space knowledge graph and available interactions, while the meta-level knowledge $\mathcal{K}=(A,G_{K},\lambda)$ comprises of its type space knowledge graph and interactions.
+Therefore, the agent’s interaction level $\mathcal{I}=(A,G_{I},\lambda)$ comprises of its instance space graph and available interactions, while the meta or knowledge level $\mathcal{K}=(A,G_{K},\lambda)$ comprises of its type space graph and meta-level actions or interactions.
 Interaction
-After the agent constructs the knowledge graphs, the meta-level recommender $\Upsilon$ recommends a candidate object $e_{i}$ and interaction type $\lambda_{p}$ for an interaction based on the agent’s knowledge graphs and meta-level policy $\pi$.
+After the agent constructs the type and instance graphs, the meta-level recommender $\Upsilon$ recommends a candidate object $e_{i}$ and interaction type $\lambda_{p}$ for an interaction based on the agent’s type and instance graphs and meta-level policy $\pi$.
 The recommender iterates through all visible objects in the agent’s instance graph $G_{I}$.
 For each object, it adds an activation relation between the agent and the object in both the instance space graph $G_{I}$ and the type space graph $G_{K}$, where ‘activating’ refers to observing or focusing on an object for potential interaction.
-For $G_{K}$, the agent’s meta-policy returns the state value and interaction probability.
+For $G_{K}$, the agent’s meta-policy returns the state value and the interaction probability.
 The state value represents the value of activating that specific object.
-Then, an object $e_{i}$ for interaction is proposed through sampling based on the values of activation of the visible objects, and interaction $\lambda_{p}$ is sampled from the interaction probability of the proposed object.
+Then, an object $e_{i}$ for interaction is proposed by sampling based on the values of activation of the visible objects, and interaction $\lambda_{p}$ is sampled from the interaction probability of the proposed object.
 Therefore, at each step, the meta-level of the agent activates a candidate object $e_{i}$ for interaction $\lambda_{p}$, setting the agent’s interaction goal $g^{i}_{p}=(e_{i},\lambda_{p})$.
@@
 eXecution
-Each agent has a single meta-level policy, whereas interaction policies are a collection of policies corresponding to different interactions.
+Each agent has a meta-level policy, whereas interaction policies are a collection of policies corresponding to different interactions.
 By following the interaction goals $g^{i}_{p}$ as directed by the meta-level, the agent interacts with objects.
@@
 Meta-level representations of type space graphs with activation relations are learned using a message-passing graph neural network.
-Nodes have abstract level attributes such as type encoding, while edges have edge type encoding for different types of relationships.
+Nodes have abstract level attributes such as type encoding, while edges have edge-type encoding for different types of relations.
 We use GAT convolutional layers Brody et al.
 (2021); Fey and Lenssen (2019) to learn graph representations given as,
-$h_{v}^{\prime}=\alpha_{v,v}\theta h_{v}+\sum_{u\in\mathcal{N}(v)}\alpha_{u,v}\theta h_{u}$
+$h_{v}^{\prime}=\alpha_{v,v}\theta h_{v}+\sum_{u\in\mathcal{N}(v)}\alpha_{u,v}\theta h_{u}$ (2)
 where $h_{v}^{\prime}$ is the updated embedded representation of node $v$, $\mathcal{N}(v)$ is the set of neighbors of node $v$, and $h_{u}$ is the representation of the neighboring node.
@@
 Meta level PPO optimizes a surrogate loss:
-$\mathcal{L}_{m}^{\text{actor}}(\theta)=-\mathbb{E}_{\mathcal{D}_{\tau_{m}}}\left[\min\left(r_{\theta}A_{m},\text{clip}(r_{\theta},1-\epsilon,1+\epsilon)A_{m}\right)\right]$
+$\mathcal{L}_{m}^{\text{actor}}(\theta)=-\mathbb{E}_{\mathcal{D}_{\tau_{m}}}\left[\min\left(r_{\theta}A_{m},\text{clip}(r_{\theta},1-\epsilon,1+\epsilon)A_{m}\right)\right]$ (3)
 with $r_{\theta}=\frac{\pi_{\theta}(.|G_{K})}{\pi_{\theta}^{r}(.|G_{K})}$ being the ratio of probabilities of actions at meta level roll-out.
 The critic loss is the mean squared error between the estimated values and the returns:
-$\mathcal{L}_{m}^{\text{critic}}(\theta)=\mathbb{E}_{\mathcal{D}_{\tau_{m}}}\left[\frac{1}{2}(V_{\pi_{\theta}}(G_{K})-R_{m})^{2}\right]$
+$\mathcal{L}_{m}^{\text{critic}}(\theta)=\mathbb{E}_{\mathcal{D}_{\tau_{m}}}\left[\frac{1}{2}(V_{\pi_{\theta}}(G_{K})-R_{m})^{2}\right]$ (4)
 Considering the mean entropy $\mathcal{H}_{\theta}$, the overall minimization objective at the meta-level becomes:
-$\mathcal{L}_{m}(\theta)=\mathcal{L}_{m}^{\text{actor}}(\theta)-\zeta_{m}^{h}\mathcal{H}_{\theta}+\zeta_{m}^{c}\mathcal{L}_{m}^{\text{critic}}(\theta)$
+$\mathcal{L}_{m}(\theta)=\mathcal{L}_{m}^{\text{actor}}(\theta)-\zeta_{m}^{h}\mathcal{H}_{\theta}+\zeta_{m}^{c}\mathcal{L}_{m}^{\text{critic}}(\theta)$ (5)
 where $\zeta_{m}^{h}$, $\zeta_{m}^{c}$ are the entropy and value coefficients, respectively.
@@
 Interaction policy learning
-Interaction policies are a repository of policies for different interactions which can be recommended by the meta-level, $\mu_{\phi}=\{\mu_{\phi_{1}},\mu_{\phi_{2}},...,\mu_{\phi_{p}}\}$ and each policy is parameterized separately.
+Interaction policies are a repository of policies for different interactions which can be recommended by the meta level, $\mu_{\phi}=\{\mu_{\phi_{1}},\mu_{\phi_{2}},...,\mu_{\phi_{p}}\}$ and each policy is parameterized separately.
 When an interaction goal with an object $e_{i}$ and interaction $\lambda_{p}$ is recommended, $g^{i}_{p}=(e_{i},\lambda_{p})$, the agent attempts to achieve it by executing a series of low-level actions from $u$.
@@
 Each interaction level PPO optimizes a surrogate loss:
-$\mathcal{L}_{p}^{\text{actor}}(\phi_{p})=-\mathbb{E}_{\mathcal{D}_{\tau_{p}}}\left[\min\left(r_{\phi_{p}}A_{p},\text{clip}(r_{\phi_{p}},1-\epsilon,1+\epsilon)A_{p}\right)\right]$
+$\mathcal{L}_{p}^{\text{actor}}(\phi_{p})=-\mathbb{E}_{\mathcal{D}_{\tau_{p}}}\left[\min\left(r_{\phi_{p}}A_{p},\text{clip}(r_{\phi_{p}},1-\epsilon,1+\epsilon)A_{p}\right)\right]$ (6)
 with $r_{\phi_{p}}=\frac{\mu_{\phi_{p}}(.|s)}{\mu_{\phi_{p}}^{r}(.|s)}$ is the ratio of probabilities of actions of policy and roll-out policy.
 The critic loss is the mean squared error between the estimated values and the returns:
-$\mathcal{L}_{p}^{\text{critic}}(\phi_{p})=\mathbb{E}_{\mathcal{D}_{\tau_{p}}}\left[\frac{1}{2}(V_{\mu_{\phi_{p}}}(s)-R_{p})^{2}\right]$
+$\mathcal{L}_{p}^{\text{critic}}(\phi_{p})=\mathbb{E}_{\mathcal{D}_{\tau_{p}}}\left[\frac{1}{2}(V_{\mu_{\phi_{p}}}(s)-R_{p})^{2}\right]$ (7)
 Considering the mean entropy $\mathcal{H}_{\phi_{p}}$, the overall minimization objective becomes:
-$\mathcal{L}_{p}(\phi_{p})=\mathcal{L}_{p}^{\text{actor}}(\phi_{p})-\zeta_{p}^{h}\mathcal{H}_{\phi_{p}}+\zeta_{p}^{c}\mathcal{L}_{p}^{\text{critic}}(\phi_{p})$
+$\mathcal{L}_{p}(\phi_{p})=\mathcal{L}_{p}^{\text{actor}}(\phi_{p})-\zeta_{p}^{h}\mathcal{H}_{\phi_{p}}+\zeta_{p}^{c}\mathcal{L}_{p}^{\text{critic}}(\phi_{p})$ (8)
 for each interaction $p$, where $\zeta_{p}^{h}$, $\zeta_{p}^{c}$ are the entropy and value coefficients, respectively.
@@
 In summary, when viewing tasks as meta-policy guided interactions with objects, problems are naturally decomposed into interaction goals.
-Unlike the models that are tightly coupled with low-level representations through end-to-end policies, in the proposed framework, agent actions are aided by meta-level knowledge guided object interactions, providing a mechanism for generalist behaviors.
-4 Results and Discussion
+Unlike the models that are tightly coupled with low-level representations through end-to-end policies, in the proposed framework, agent actions are aided by the interpretation of meta-level knowledge to guide object interactions, providing a mechanism for generalist behaviors.
+Results and discussion
 Experiments
@@
 To assess the capabilities of the proposed agents, we need a challenging environment having ingredients for higher-order reasoning to solve tasks.
-MiniGrid world environments Chevalier-Boisvert et al.
+MiniGrid world navigation environments Chevalier-Boisvert et al.
 (2018) are suitable for this purpose; they are procedurally generated environments where the agent interacts with objects such as doors, keys, balls, and boxes.
-ObstructedMaze-Full
-ObstructedMaze-1Dlhb
-Figure 3: Environments: ObstructedMaze-Full is used to experimentally evaluate the agent’s performance on the designed tasks, while ObstructedMaze-1Dlhb is used to evaluate the transferability of interaction policies.
 The obstructed maze environments, shown in Figure 3, are particularly challenging due to partial observability, stochasticity, and sparse rewards, with only a few claims for solving it.
@@
 Keys are hidden inside boxes, but the agent cannot see inside the boxes without toggling.
-The default task is referred to as Task 0.
-Furthermore, we identify the following new task scenarios, which were presented to the agent during testing but not during training:
+(a) ObstructedMaze-Full
+(b) ObstructedMaze-1Dlhb
+Figure 3: Environments: ObstructedMaze-Full is used to experimentally evaluate the agent’s performance on the designed tasks, while ObstructedMaze-1Dlhb is used to evaluate the transferability of interaction policies.
+The default task is referred to as Task 0. Furthermore, we identify the following new task scenarios, which were presented to the agent during testing but not during training:
   • No context in training (Task 1) - a relation that the agent did not come across during training.
@@
 It is only done once, not repeatedly.
-Agents learning to interact using type space knowledge
+Learning to interact using type space knowledge
+We devised two variants: KIX-A and KIX-R.
+Both variants learn a meta-policy and a collection of interaction policies.
+While the KIX-A variant combines both reaching objects and interacting with them, the KIX-R variant learns an additional interaction policy referred to as the ‘reachability’ policy, which is used to reach objects prior to interacting with them.
+All policies, including meta and interactions, are implemented as actor-critic PPO policies.
 initialize model weights $\theta,\phi$ , parameters, $T$
 for  $t=1$ to $T$ do
-       /* collect exp for meta exp size K */
+   /* collect meta experiences */
  for $k=1$ to $K$ do
@@
 Algorithm 1 KIX procedure
-We devised two variants: KIX-A and KIX-R.
-Both variants learn meta-policy and a collection of interaction policies.
-While the KIX-A variant combines both reaching objects and interacting with them, the KIX-R variant learns an additional interaction policy referred to as the ‘reachability’ policy, which is used to reach objects prior to interacting with them.
-All policies, including meta and interactions, are implemented as actor-critic PPO policies.
 Algorithm 1 provides an overview of the procedure.
@@
 Edges in a type space graph have the same meaning as those in an instance space graph and are encoded as one-hot encoded edge attribute vectors.
-Figure 4: Agents learning to interact: The learning progress of the agents KIX-A (blue) and KIX-R (red) is shown as returns over steps, which illustrate both the meta-level and interactions such as ‘open’, ‘open with key’, ‘reveal’, ‘pickup’ (not shown for KIX-R), ‘drop’, and ‘reach’ in the ObstructedMaze-Full environment.
 Agents learn how to perform low-level actions in the execution space using interaction policies.
@@
 At the meta-level, it is rewarded with 0.1 for a successful proposed interaction and $1+r$ for achieving the environment goal.
-The base agent, BASE, has a similar actor-critic architecture and receives environment rewards.
+The base agent, BASE, is an end to end policy learner with a similar actor-critic architecture and receives environment rewards.
 The training hyper-parameters are given in Table 1.
 Learning rates were determined using grid search with values 1e-3, 1e-4, 1e-5, and 1e-6.
-All the agents are trained for $\sim 6.2\times 10^{8}$ steps, using 32 parallel workers.
-As shown in Figure 4, as training progresses, agents learn to interact with objects and associate higher-level structured knowledge with these interactions using object types and knowledge graphs, as indicated by higher meta-level returns.
+All the agents are trained for $\sim 6.2\times 10^{8}$ steps, using 32 parallel workers in a high performance computing cluster.
+Figure 4: Agents learning to interact: The learning progress of the agents KIX-A (blue) and KIX-R (red) is shown as returns over steps, which illustrate learning both the meta-level and interactions such as ‘reach’ (only applicable for KIX-R), ‘open’, ‘open with key’, ‘reveal’, ‘pickup’, ‘drop’ (not shown) in the ObstructedMaze-Full environment.
+As shown in Figure 4, as training progresses, agents learn to interact with objects and associate higher-level structured knowledge with these interactions using object types and type and instance space graphs, as indicated by higher meta-level returns.
 This, in turn, leads to better interaction experiences and further improvements in interaction policies.
 Furthermore, the KIX-R variant first focuses on learning how to reach different objects, but once it learns to reach, it begins to learn interactions in fewer steps and with much higher returns compared to the KIX-A variant.
-It suggests that the agents learn higher-level relations and apply these learned relations and interactions with objects to solve problems.
-Meta-level knowledge-guided interactions aid agents
+It suggests that the agents learn higher-order relations and apply these learned relations and interactions with objects to solve problems.
+Meta-level guided interactions aid agents
 Our approach is motivated by the idea that meta-level guided interactions with objects help agents in generalizing while performing tasks, rather than acting directly at the execution level.
@@
 (a) Meta training
-(b) Evaluation task returns
+(b) Evaluation of task returns
 Figure 6: Evaluation return profiles show that interaction policies can be transferred from one environment to another.
@@
 In this environment, the agent’s goal is to pick up a blue ball hidden in a two-room maze.
-The rooms are separated by a door, which is obstructed by a ball and locked.
+The rooms are separated by a door that is obstructed by a ball and locked.
 The keys are hidden inside boxes.
 Agents employ previously acquired interaction policies from the base task in the ObstructedMaze-Full environment while only learning the meta-policies.
-During evaluation, agents use the learned meta-policy, but the interaction policies come from the ObstructedMaze-Full environment itself.
+During evaluation, agents use the learned meta-policy, but the interaction policies are from the ObstructedMaze-Full environment itself.
 Figure 6 shows the meta training returns and evaluation return profiles, similar to the main environment, for both agent variants.
 It is evident that agents have gained the ability to exploit interaction policies from the previous environment, and the second variant, which employs a separate reachability policy, appears to have benefited more.
-5 Conclusion
+Conclusion
 In this work, we proposed a metacognitive framework that promotes generalist behaviors through three main contributions.
 First, generalist behaviors necessitate abstraction and composition; however, current models lack natural abstractions and compositions that can preserve both object abstractions and relational knowledge structures.
-In our model, type space knowledge graphs are used to represent and ground natural abstractions and compositions while also preserving relational knowledge between objects.
-Second, rather than directly acting in the environment, interactions with objects are critical in performing the tasks.
-To accomplish this, introduce an intermediary interaction level.
-The agent’s meta-level learns from type space knowledge graphs and makes recommendations on which objects to interact with and how to interact with them.
-In this sense, our model functions as an interaction goal-generating machine.
+In our model, type-space graphs are used to represent and ground natural abstractions and compositions while also preserving relational knowledge between objects.
+Second, rather than acting directly in the environment, interactions with objects are critical in performing tasks.
+To accomplish this, we introduced an intermediary interaction level.
+The agent’s meta-level interprets abstract relational knowledge, learns a meta-policy using type-space graphs, and makes recommendations on which objects to interact with and how to interact with them.
+In this sense, our model functions as an interpretation-based interaction goal-generating machine.
 Third, interactions employing interaction policies enable the necessary loose coupling required for generalization by acting as intermediaries.
 We demonstrated that metacognitive knowledge-guided interactions aid agents in generalization.
-Lastly, Our framework is modular and adaptable, allowing for the integration of knowledge structures and control variables at both the meta and interaction levels.
+Lastly, Our framework is modular and adaptable, enabling the integration of knowledge structures and control variables at both the meta and interaction levels.
 This framework is promising, particularly for domains where higher-order knowledge and interaction compositions can be exploited, such as robotics, autonomous and cognitive systems.

### v4

@@
-KIX: A Knowledge and Interaction-Centric Metacognitive Framework for Task Generalization
+Knowledge-Centric Metacognitive Learning
 Abstract
-People aptly exhibit general intelligence behaviors through flexible problem-solving and the ability to adapt to novel situations by reusing and applying high-level knowledge acquired over time.
-In contrast, artificial agents tend to be specialists, lacking such generalist behaviors.
-To bridge this gap, artificial agents will require understanding and exploiting critical structured knowledge representations.
-We introduce a metacognitive reasoning framework, Knowledge-Interaction-eXecution (KIX), and argue that interactions with objects, by leveraging a type space, facilitate the learning of transferable interaction concepts and promote generalization.
-This framework offers a principled approach for integrating knowledge into reinforcement learning and holds promise as an enabler for generalist behaviors in artificial intelligence, robotics, and autonomous systems.
+Interactions are central to intelligent reasoning and learning abilities, with the interpretation of abstract knowledge guiding meaningful interaction with objects in the environment.
+While humans readily adapt to novel situations by leveraging abstract knowledge acquired over time, artificial intelligence systems lack principled mechanisms for incorporating abstract knowledge into learning, leading to fundamental challenges in the emergence of intelligent and adaptive behavior.
+To address this gap, we introduce knowledge-centric metacognitive learning based on three key principles: natural abstractions, knowledge-guided interactions through interpretation, and the composition of interactions for problem solving.
+Knowledge learning facilitates the acquisition of abstract knowledge and the association of interactions with knowledge, while object interactions guided by abstract knowledge enable the learning of transferable interaction concepts, abstract reasoning, and generalization.
+This metacognitive mechanism provides a principled approach for integrating knowledge into reinforcement learning and offers a promising pathway toward intelligent and adaptive behavior in artificial intelligence, robotics, and autonomous systems.
 Introduction
-The ability to actively reason, flexibly evaluate options and learn from experience is a hallmark of human intelligence.
-Core to this intelligence is the human ability to identify natural decompositions within complex environments and tasks with ease, enabling flexible and generalist behaviors.
-Replicating such autonomy and general intelligence abilities in artificial agents continues to pose significant challenges, which we argue necessitate a knowledge and interaction-centric metacognitive approach.
-Meta-reasoning about next actions, their values, and possible next states underpins flexible decision-making in novel situations, problem solving, concept learning, and knowledge acquisition.
-This form of reasoning extends beyond mere memory lookup.
-More broadly, meta-reasoning functions as a sense-making process within a knowledge system, enabling people to intuitively direct their learning to discern relations between entities, interact with the same entities in diverse ways, and ground these interactions for both short- and long-term goals.
-Metacognitive reasoning thus emerges as a foundational mechanism for generalization.
-In contrast, meta-reasoning systems in artificial intelligence are primarily focused on resource allocation problems.
-While resource management is essential for efficient functioning, it is crucial to ideate metacognitive generalization, illustrated in Figure 1, to enable generalist behaviors in artificial agents.
+A hallmark of human intelligence is the capacity to construct and interpret abstract knowledge.
+Interpretation grounded in such abstractions enables meaningful interaction with the environment, forming the foundation for knowledge learning, flexible planning, and decision-making.
+It is fascinating how effortlessly people discern natural abstractions and leverage them to interact with their environments in flexible and adaptive ways.
+In contrast, contemporary artificial intelligence systems lack the principled integration of abstract knowledge into learning, resulting in fundamental limitations in the emergence of intelligent behavior.
+For example, autonomous systems trained solely on object-level data often fail to respond to changes in world characteristics due to the absence of integrated knowledge learning.
+Such gaps in knowledge learning lead to non-robust, unreliable, and specialized behavior in unknown and complex scenarios.
+Machine learning methods Goodfellow et al.
+(2016); Bishop and Nasrabadi (2006) learn a mapping function by optimizing an objective between predictions and targets; however, these targets lack the underlying know-how required for world understanding.
+In reinforcement learning Sutton and Barto (2018), agents collect experiences through interaction with their environments and learn policies for action selection, yet even model-based approaches, despite learning a dynamics model, fail to integrate abstract knowledge into learning and world models.
+Learning, at its core, is knowledge learning for an abstract understanding of the world and responding adaptively through the use of such knowledge.
+Addressing these gaps necessitates a knowledge-centric metacognitive learning approach.
+Meta-reasoning about subsequent actions underpins flexible decision-making.
+More broadly, meta-reasoning acts as a sense-making process within knowledge systems, enabling people to intuitively direct their learning to discern abstract relational structure and ground object interactions.
+However, meta-reasoning in artificial systems predominantly acts as a resource controller, monitoring lower-level resources and managing resource allocation for efficient functioning.
+While such resource management is essential, a resource-controller–only view is severely limiting for facilitating abstract reasoning and adaptive behaviors.
+Meta-reasoning must pivot toward interpretation-based reasoning within abstract knowledge.
+This paradigm of abstract interpretive reasoning and learning – termed metacognitive reasoning and learning – is central to much of knowledge learning, interaction with objects in the physical world, and manifested in the emergence of autonomous intelligent behaviors.
+Figure 1 illustrates this process in a navigation domain.
+To develop an understanding of the world, an agent must learn abstract knowledge of objects where object types serve as natural abstractions, their relations, and the ways in which they can be interacted with.
+Interpretation over this abstract knowledge guides interactions enabling flexible problem solving.
 (a) Overview
 (b) Approach
-Figure 1: a) Illustration of metacognitive reasoning process in a navigation domain, where an agent (depicted as a red triangle) must interpret its environment to achieve the goal of retrieving a blue ball located behind a locked door and various obstacles.
-This process requires reasoning over abstract object types and their relations, which in turn guide the agent’s interactions with objects.
-Such abstract reasoning fundamentally relies on higher-order abstract structures and compositional interactions over these abstractions.
-b) Learning generalist behaviors requires developing a natural and abstract understanding of the environment, as well as meta-reasoning and interaction abilities.
-The metacognitive knowledge-guided approach exploits inherent abstractions and relations between objects to generate interaction goals, enabling the agent to interact with objects and generalize to solve complex tasks.
-Metacognitive reasoning is fundamentally grounded in interaction, requiring agents to draw upon their interpretations of abstract knowledge to guide their interactions.
-Interpretation precedes and shapes interaction; it involves developing both natural and abstract understandings of the environment that can be deployed as needed to make sense of the current state and plan subsequent interactions.
-To illustrate this process, consider the example in Figure 1a, which depicts an agent and its interactions in a navigation domain.
-To build knowledge of the environment, an agent must develop a conceptual understanding of the objects in the environment and how to interact with them.
-As some objects may even appear similar, the agent must decipher their abstract types.
-Learning to interact then involves learning which interactions are possible with these abstract object types and the relations between them.
-Type serves as a natural abstraction, and interpretation thus incorporates this meta-level reasoning over abstract knowledge.
-The utility of interactions is then learned through meta-level policies in this abstract space, which can subsequently be leveraged to recommend interactions and execute compositions of interactions at a lower level, thereby facilitating problem-solving.
-Our proposal is based on three main principles for flexibly solving complex tasks.
-Natural abstraction: Real-world environments are naturally composed of entities categorized by their types, similar to a natural hierarchy of objects and classes.
-Entities in the instance space inherit properties and relations from a higher-level type space.
-Type-space graphs represent the underlying objects and their relations in environments.
-This abstraction is where the knowledge resides and it is the knowledge or meta level.
-Interactions: An agent’s metacognitive process employs the type-space graphs to create a plan specifying which entities to interact with and the nature of these interactions to complete tasks and achieve goals.
-While the metacognitive policy guides agents based on abstract knowledge, agents interact with entities in the instance space which are instantiations of abstract types.
-This is the interaction level.
-Composition of interactions: The agent interacts with meta-guided objects as a recommended interaction type and executes low-level policies corresponding to a guided interaction type.
-Therefore, a composition of these interactions as interaction pathways for graph traversal at both the meta level and the instance level facilitates execution of low-level actions which are loosely coupled with abstract knowledge through interactions.
-Integrating these principles into metacognitive learning, reasoning, and control provides the levers and connecting points for instilling generalist behaviors in agents.
-Agents’ flexible behaviors stem from their ability to interpret knowledge, learn how to interact with objects through meta level policies and leverage interaction pathways enabling quick adaptation.
-The proposed metacognitive framework captures the essence of generalist behaviors and may have far-reaching implications across various domains.
-We describe the key components of the framework that exemplify generalist characteristics, followed by an experimental evaluation in a navigation domain and discuss how we envision it enabling flexibility through meta-level guided interactions.
+Figure 1: a) Illustration of metacognitive approach in a navigation domain, where an agent must develop an abstract understanding of its environment and interpret it to achieve the goal of retrieving a blue ball located behind a locked door, blocked by obstacles, with the key hidden inside a box.
+b) A knowledge centric metacognitive learning approach exploits inherent object abstractions and relations to generate interaction goals and compose interactions, thereby enabling the emergence of intelligent and flexible behavior.
+Our proposal is based on three main principles for knowledge-centric learning and problem-solving.
+Natural abstraction: Real-world environments are naturally composed of entities categorized by their types, similar to an inherent hierarchy of objects and classes.
+Objects in the instance space inherit properties and relations from a higher-level type space.
+Type space graphs provide an abstract representation of objects and their relationships within the environment.
+The abstraction space is where the knowledge resides and is referred to as the knowledge or meta level.
+Interactions: An agent’s metacognitive process employs type space graphs to create plans that specify which entities to interact with and the nature of these interactions in order to achieve its goals.
+While the meta-level policy guides the agent based on abstract knowledge, the agent executes interactions in the instance space, where objects are type instantiations.
+This constitutes the interaction level.
+Composition of interactions: The meta-level guides the agent to interact with recommended objects using interaction types, and the agent executes low-level policies corresponding to guided interactions.
+The composition of these interactions forms interaction pathways for graph traversal at both the meta and instance level, thereby facilitating the execution of low-level actions.
+The process is loosely coupled through interactions, enabling flexible and adaptive behavior.
+Integrating these principles into metacognitive learning provides foundational mechanisms for endowing agents with knowledge learning and adaptive capabilities, arising from their ability to interpret abstract knowledge, learn object interactions guided by meta-level policies, and leverage interaction pathways for adaptation.
+The proposed metacognitive learning approach captures the essence of abstract reasoning, knowledge learning, and flexible behavior, with potentially far-reaching implications across domains that require abstract reasoning and interaction with the physical world.
+In the remaining of this article, we describe the key components and exemplifying characteristics of the approach, followed by an experimental evaluation in a navigation domain, and discuss how it enables adaptive problem solving through knowledge–guided interaction learning.
 Background
@@
 Related Work
-Compositionality and abstractions are central to generalization learning, with abstraction requiring a mechanism for grounding representations in actual objects.
+Compositionality and abstractions are central to generalization, with abstraction requiring a mechanism for grounding representations in objects.
 An object-centric model Veerapaneni et al.
-(2020) was presented in which object representations were constructed from raw observations and simpler local functions with uniformity, which were then used for prediction and planning.
-Similarly, a symbolic object-centric representation Vijay et al.
+(2020) was presented in which object representations were constructed from raw observations and local functions with uniformity were then used for prediction and planning.
+A symbolic object-centric representation Vijay et al.
 (2019) was used to study the impact of adding prior knowledge through relations in an object graph.
 These studies focus primarily on object space, which inhibits the abstraction of objects and relational structures naturally occurring in the real world and used for decision-making.
-Key questions include what a natural abstraction should be and how it can be used to compose interactions as building blocks leveraging natural relational structures.
+Key questions include what a natural abstraction should be and how it can be used to compose interactions as building blocks for problem solving.
 We address these largely overlooked questions by preserving abstract knowledge and relational higher-order structures in the form of knowledge graphs, where abstractions are grounded using type space.
@@
 (2020).
-These methods enable agents to learn hierarchy of policies, but they remain in the object space for both low- and high-level actions and do not meet a key meta-reasoning demand: type abstractions and loose coupling between abstraction and execution space.
+These methods enable agents to learn hierarchy of policies, but they remain in the object space for both low- and high-level actions and do not meet key meta-reasoning demands: type abstractions and loose coupling.
 Methods such as Ten et al.
@@
 (2005) develop internal goals for empowerment, focusing on aspects different than ours.
-On the other hand, neuro-symbolic reasoning approaches Hitzler and Sarker (2022); Zhang and Sheng (2024) aim to overcome the limitations of purely symbolic or purely neural methods by integrating them into hybrid models that combine the learning capacity of neural networks with the logical inference and interpretability of symbolic systems.
-However, significant challenges persist in facilitating interaction and integration between these disparate paradigms to ensure a seamless, flexible, and semantically grounded natural unification.
-Our approach meets these demands by leveraging natural higher-order knowledge structures in the form of type space graphs for meta-level policy and facilitating interactions with objects via an intermediary between meta-level and execution space, thus enabling loose coupling and reuse of low-level policies.
-The appeal of our approach is that interaction policies can be swapped in and out depending on the interaction type.
-Our work distinguishes itself by incorporating relational abstract knowledge structures to guide interactions that govern low-level execution.
+Goal-conditioned reinforcement learning Liu et al.
+(2022); Pong et al.
+(2018) seeks to support multi-task generalization by conditioning policies on goals, yet remains challenged due to the lack of integrated knowledge based reasoning and learning in complex tasks.
+We address these limitations through a metacognitive approach, in which abstract knowledge constrains learning, generates and guides interactions, enabling flexible behavior.
+Neuro-symbolic approaches Hitzler and Sarker (2022); Zhang and Sheng (2024) aim to overcome the limitations of purely symbolic or neural methods by integrating them into hybrid models that combine the learning capacity of neural networks with the logical inference and interpretability of symbolic systems.
+However, significant challenges persist in facilitating interaction and integration between these disparate paradigms to ensure a seamless and semantically grounded unification.
+Our approach addresses these demands by leveraging natural higher-order knowledge structures, represented as type-space graphs, for meta-level policy learning, and by mediating object interactions via an intermediary between the meta and execution levels, thus enabling loose coupling and reuse of low-level policies and interaction policies can be swapped in and out based on interaction type.
+Our work is distinguished by its principled incorporation of relational abstract knowledge to guide interactions that govern low-level execution.
 Meta-reasoning
 Meta-reasoning is often defined as the set of processes that monitor our thought processes and regulate resource allocation Ackerman and Thompson (2017).
-Its basic structure Nelson (1990) consists of two levels: meta and object.
+Its basic structure Nelson (1990) consists of levels: meta and object.
 The meta level’s functions include control and monitoring.
@@
 The monitoring process allows the meta controller to track the status of object-level processes and actions, as well as the consequences of those actions.
-However, current meta-reasoning, primarily as a resource manager, is insufficient for exhibiting generalist behaviors, necessitating generalist meta-cognitive approaches such as the one proposed here.
+However, current meta-reasoning, primarily as a resource manager, is insufficient for exhibiting intelligent and generalist behaviors, necessitating a meta-cognitive approach.
 Graph Neural Networks
@@
 One typical task is to predict the label or value of the entire graph.
-Graph neural networks (GNN) (Gilmer et al.
+Graph neural networks (GNN), Gilmer et al.
 (2017); Hamilton et al.
-(2017) and others) use node connectivity as well as node and edge features to learn a permutation-invariant embedding vector for each node $h_{v},v\in\mathcal{V}$ and the entire graph $h_{\mathcal{G}}$.
+(2017) and others, use node connectivity as well as node and edge features to learn a permutation-invariant embedding vector for each node $h_{v},v\in\mathcal{V}$ and the entire graph $h_{\mathcal{G}}$.
 Typically, the representation is learned through message passing or neighborhood feature aggregation, in which node embedding vectors are updated in each layer based on message aggregation from neighbors:
 $h_{v}^{(k)}=\mathcal{F}^{\prime}\left(h_{v}^{(k-1)},\bigoplus_{u\in\mathcal{N}(v)}\mathcal{F}^{\prime\prime}\left(h_{v}^{(k-1)},h_{u}^{(k-1)},\mathtt{e}_{u,v}\right)\right)$ (1)
-where $h_{v}^{(k)}$ is the updated representation of node $v$ embedding in layer $k$, $\mathcal{N}(v)$ is the set of neighboring nodes of node $v$, $h_{u}^{(k-1)}$ is the representation of neighboring node $u$ in layer $k-1$, $\bigoplus$ is a differentiable permutation invariant aggregation function, $\mathcal{F}^{\prime}$ and $\mathcal{F}^{\prime\prime}$ are differentiable functions usually neural networks.
+where $h_{v}^{(k)}$ is the updated representation of node $v$ embedding in layer $k$, $\mathcal{N}(v)$ is the set of neighboring nodes of node $v$, $h_{u}^{(k-1)}$ is the representation of neighboring node $u$ in layer k1k-1, $\bigoplus$ is a differentiable permutation invariant aggregation function, $\mathcal{F}^{\prime}$ and $\mathcal{F}^{\prime\prime}$ are differentiable functions usually neural networks.
 Node features are transformed prior to aggregation, and the methods used to transform and aggregate features differ between graph network variants.
@@
 To reduce the variance in the expected return estimate, a baseline is subtracted from it; typically, a learned value function is used as the baseline.
-The resulting quantity $A(s,a)=Q(s,a)-V(s)$ can be viewed as an estimate of the advantage of action $a$ in state $s$.
+The resulting quantity A(s,a)=Q(s,a)V(s)A(s,a)=Q(s,a)-V(s) can be viewed as an estimate of the advantage of action $a$ in state $s$.
 This method is called advantage actor critic (A2C) Sutton and Barto (2018); Degris et al.
@@
 Several improved methods have been developed to address sample efficiency, stability, and training difficulty.
-Proximal policy optimization (PPO) Schulman et al.
-(2017) is an on-policy policy optimization technique that addresses concerns in actor-critic methods using clipped objectives.
-We adapt PPO for learning policies.
-Metacognitive framework
-Our proposal for metacognitive generalization centers around the ideas of natural decomposition of tasks that an agent can accomplish through interpretation of knowledge and interpretation guided interactions with objects.
-The meta-policy recommends localized agent-object interactions, which the agent executes following interaction policies to construct a compositional solution.
-Here, we discuss the main ideas of the framework (see Figure 2) and argue that it exhibits task-generalization characteristics.
-Figure 2: The meta-cognitive generalization framework consists of three levels: knowledge ($\mathcal{K}$), interaction ($\mathcal{I}$), and execution ($\mathcal{X}$), which are based on abstraction and computation for providing generalization capabilities.
-The knowledge level contains type space knowledge graphs, which represent higher order abstractions.
-The instance space contains object instance graphs.
-The interaction level serves as an intermediary, making recommendations for interactions.
-The execution level is the lowest level at which the agent acts based on the recommended interaction subgoals.
-Computations proceed through recognition, recommendation, execution, and interactions to collect experiences for updates to interaction and meta policies.
+Policies are learned by adapting Proximal policy optimization (PPO) Schulman et al.
+(2017), an on-policy policy optimization that addresses concerns in actor-critic methods using clipped objectives.
+Metacognitive Learning Framework
+Our proposal for metacognitive learning centers on the interpretation of abstract knowledge to guide interactions with objects.
+A meta-policy recommends localized agent–object interactions, which the agent executes via interaction policies to construct a compositional task solution.
+Here, we outline the main ideas of the metacognitive learning framework, Knowledge-Interaction-eXecution (KIX) (see Figure 2).
+Figure 2: The metacognitive learning framework comprises three levels: knowledge or meta ($\mathcal{K}$), interaction ($\mathcal{I}$), and execution ($\mathcal{X}$), based on abstraction and computation.
+The knowledge level contains type space knowledge graphs that represent higher-order abstractions, while the instance space comprises object instance graphs.
+The interaction level serves as an intermediary, making recommendations for interactions that the agent executes at the execution level.
+Computation proceeds through cycles of recognition, recommendation, execution, and interaction, to collect experiences for updates to both the meta-level and interaction policies.
 Knowledge
@@
 The agent can perform low-level actions $u=\{u_{1},u_{2},$...\} within the environment.
-When making decisions, the agent organizes its perceived knowledge into graphs: an instance-level graph and a type-level graph.
-An instance-level graph is an ordered pair $G_{I}=(e,\Lambda)$ of observed objects $e$ and their relations $\Lambda=\{\Lambda_{1},\Lambda_{2},...,\Lambda_{\kappa}\}$.
-A type-level graph is an ordered pair $G_{K}=(E,\Lambda)$ of entity types $E$ and their type level or abstract relations $\Lambda$.
+When making decisions, the agent organizes its knowledge into instance, and type graphs.
+An instance graph is an ordered pair $G_{I}=(e,\Lambda)$ of observed objects $e$ and their relations $\Lambda=\{\Lambda_{1},\Lambda_{2},...,\Lambda_{\kappa}\}$.
+A type graph is an ordered pair $G_{K}=(E,\Lambda)$ of entity types $E$ and their abstract relations $\Lambda$.
 Relations, whether between objects or entity types, retain the same meaning, such as adjacency indicating that two objects are close to one another.
-The function $\Phi:X\to G_{I}$ maps the agent’s partially observed view to the instance graph by capturing observed objects and deriving their relationships.
-The function $\Psi:G_{I}\to G_{K}$ maps instance space graphs to type space graphs, a direct translation since the agent’s recognition system contains types corresponding to the objects.
-If the agent encounters a novel object without a corresponding type, it must create a new type in its type space or assign the object to a related type.
-In the framework, $\Psi$ is learnable and represents a belief in mapping objects to types.
+The function $\Phi:X\to G_{I}$ maps the agent’s partially observed view to the instance graph by capturing observed objects and relations.
+The function $\Psi:G_{I}\to G_{K}$ maps instance graphs to type graphs, a direct translation since the agent’s recognition system contains types corresponding to the objects.
+If the agent encounters a novel object, it must create a new type in its type space or assign the object to a related type.
+In the framework, $\Psi$ is learnable and represents a belief, mapping objects to types.
 However, this article treats the mapping as fixed, from object nodes in the instance graph $G_{I}$ to type nodes in the type graph $G_{K}$.
-As the agent interacts with objects, the environment changes, and its recognition system captures a new partially observed view, resulting in new instance and type space graphs.
+As the agent interacts with objects, the environment changes, and its recognition system captures a new partially observed view, resulting in new instance and type graphs.
 The agent’s knowledge level is also referred to as the meta-level.
-At the meta-level, the agent uses a meta-policy that accepts a type space graph and returns interactions or meta-level actions for object interaction.
+At the meta-level, the agent uses a meta-policy that accepts a type graph and returns interactions or meta-level actions for object interaction.
 Interactions are represented as $\lambda=\{\lambda_{1},\lambda_{2},...,\lambda_{p}\}$.
-Similarly, for each interaction type, the agent uses an interaction policy to guide low-level actions $u$ within the environment.
-Therefore, the agent’s interaction level $\mathcal{I}=(A,G_{I},\lambda)$ comprises of its instance space graph and available interactions, while the meta or knowledge level $\mathcal{K}=(A,G_{K},\lambda)$ comprises of its type space graph and meta-level actions or interactions.
+For each interaction type, the agent uses an interaction policy to execute low-level actions $u$ within the environment.
+Therefore, the agent’s interaction level $\mathcal{I}=(A,G_{I},\lambda)$ comprises its instance graph and available interactions, while the knowledge or meta level $\mathcal{K}=(A,G_{K},\lambda)$ comprises its type graph and meta level actions or interactions.
 Interaction
@@
 The recommender iterates through all visible objects in the agent’s instance graph $G_{I}$.
-For each object, it adds an activation relation between the agent and the object in both the instance space graph $G_{I}$ and the type space graph $G_{K}$, where ‘activating’ refers to observing or focusing on an object for potential interaction.
+For each object, it adds an activation relation between the agent and the object in both the instance graph $G_{I}$ and the type graph $G_{K}$, where ‘activating’ refers to observing or focusing on an object for potential interaction.
 For $G_{K}$, the agent’s meta-policy returns the state value and the interaction probability.
@@
 Meta-policy Learning
-Meta-level representations of type space graphs with activation relations are learned using a message-passing graph neural network.
+Meta-level representations of type graphs with activation relations are learned using a message-passing graph neural network.
 Nodes have abstract level attributes such as type encoding, while edges have edge-type encoding for different types of relations.
@@
 In summary, when viewing tasks as meta-policy guided interactions with objects, problems are naturally decomposed into interaction goals.
-Unlike the models that are tightly coupled with low-level representations through end-to-end policies, in the proposed framework, agent actions are aided by the interpretation of meta-level knowledge to guide object interactions, providing a mechanism for generalist behaviors.
+Unlike the models that are tightly coupled with low-level representations through end-to-end policies, in the proposed framework, agent actions are aided by the interpretation of meta-level knowledge to guide object interactions, providing a metacognitive mechanism for knowledge learning, abstract reasoning and generalist behaviors.
 Results and Discussion
 Experiments
-Our goal is to examine whether meta-level knowledge-guided interactions benefit agents in task generalization.
-To assess the capabilities of the proposed agents, we need a challenging environment having ingredients for higher-order reasoning to solve tasks.
+We aim to examine whether agents can learn to associate abstract relational knowledge with object and interaction types, leverage this knowledge for interaction learning and guide interactions to exhibit flexible behavior when solving complex tasks.
+To assess the proposed agents, we need a challenging environment having ingredients for higher-order reasoning to solve tasks.
 MiniGrid world navigation environments Chevalier-Boisvert et al.
@@
 (b) ObstructedMaze-1Dlhb
-Figure 3: Environments: ObstructedMaze-Full is used to experimentally evaluate the agent’s performance on the designed tasks, while ObstructedMaze-1Dlhb is used to evaluate the transferability of interaction policies.
+Figure 3: Environments: ObstructedMaze-Full is used to experimentally evaluate the agent’s knowledge learning and flexible task solving capabilities, while ObstructedMaze-1Dlhb is used to evaluate the transferability of interaction policies.
 The default task is referred to as Task 0. Furthermore, we identify the following new task scenarios, which were presented to the agent during testing but not during training:
-  • No context in training (Task 1) - a relation that the agent did not come across during training.
+  • No context in training (Task 1) – a relation that the agent did not come across during training.
 It is simulated by placing the goal inside an empty box.
-During training, the blue ball was never in a box, although the agent encountered boxes and balls during training, but boxes contained keys.
-  • Different context in testing (Task 2) - using a relation that the agent encountered during training but in a new context during testing.
+During training, the blue ball was never in a box, although the agent encountered boxes and balls during training, but the boxes contained keys.
+  • Different context in testing (Task 2) – using a relation that the agent encountered during training but in a new context during testing.
 It is simulated by locking doors in the central room and placing keys directly in the room.
-During training, doors in the central room were never locked, nor was there direct access to keys.
-  • Dynamic changes in the environment (Task 3) - when some event occurs, the environment changes.
-During training, the goal location remained unchanged.
-To simulate, as soon as the agent enters the room with the goal, relocate the goal to one of the neighboring rooms.
-It is only done once, not repeatedly.
-Learning to interact using type space knowledge
+During training, the doors in the central room were never locked, nor was there direct access to the keys.
+  • Dynamic changes in the environment (Task 3) – when some event occurs, the environment changes.
+During training, the goal location in an episode remained unchanged.
+To simulate this, as soon as the agent enters the room with the goal, relocate the goal to one of the neighboring rooms.
+It is done only once, not repeatedly.
+Knowledge Learning to Interact
 We devised two variants: KIX-A and KIX-R.
@@
 Algorithm 1 provides an overview of the procedure.
-The agent’s partially observed view $X$ contains objects with identifiers and attributes like color and type.
+The agent’s partially observed view $X$ contains objects with identifiers and attributes such as color and type.
 The function $\Phi$ takes $X$ and creates an instance space graph by using visible objects as nodes, including the agent, and determining relationships between nodes such as ‘visible’, ‘adjacent’, and ‘carrying’.
@@
 The meta-level actor and critic are both fully connected networks, with the critic learning a single value and the actor learning the likelihood of taking the meta-level actions ‘pickup’, ‘drop’, ‘reveal’, ‘open’, or ‘open with key’, each of which corresponds to an interaction policy.
+Parameter Value Parameter Value Parameter Value
+meta exp size 8 learning rate 1e-05 gamma 0.99
+exp size 64 entropy coeff 1e-04 gae lambda 0.95
+meta minibatch size 64 value coeff 1.0 ppo clip 0.2
+minibatch size 64 meta entropy coeff 1e-04 ppo steps 4
+meta learning rate 1e-05 meta value coeff 1.0
 Table 1: Training hyper-parameters
-Parameter Value Parameter Value
-meta exp size 8 gae lambda 0.95
-exp size 64 entropy coeff 1e-04
-meta minibatch size 64 value coeff 1.0
-minibatch size 64 meta entropy coeff 1e-04
-meta learning rate 1e-05 meta value coeff 1.0
-learning rate 1e-05 ppo clip 0.2
-gamma 0.99 ppo steps 4
 The environment rewards the agents with $r=1-(c/c_{T})$ where $c$ is the number of steps taken and $c_{T}$ is the max possible steps.
@@
 All the agents are trained for $\sim 6.2\times 10^{8}$ steps, using 32 parallel workers in a high performance computing cluster.
-Figure 4: Agents learning to interact: The learning progress of the agents KIX-A (blue) and KIX-R (red) is shown as returns over steps, which illustrate learning both the meta-level and interactions such as ‘reach’ (only applicable for KIX-R), ‘open’, ‘open with key’, ‘reveal’, ‘pickup’, ‘drop’ (not shown) in the ObstructedMaze-Full environment.
-As shown in Figure 4, as training progresses, agents learn to interact with objects and associate higher-level structured knowledge with these interactions using object types and type and instance space graphs, as indicated by higher meta-level returns.
+Figure 4: Agents learning to interact : The learning progress of both the agents KIX-A (blue) and KIX-R (red) is shown as returns over steps, demonstrating the learning of meta-level as well as interaction policies such as ‘reach’ (only applicable for KIX-R), ‘open’, ‘open with key’, ‘reveal’, ‘pickup’, ‘drop’ (not shown) in the ObstructedMaze-Full environment.
+As shown in Figure 4, as training progresses, agents learn to interact with objects and associate higher-level structured knowledge with these interactions using object types and type and instance graphs, as indicated by higher meta-level returns.
 This, in turn, leads to better interaction experiences and further improvements in interaction policies.
 Furthermore, the KIX-R variant first focuses on learning how to reach different objects, but once it learns to reach, it begins to learn interactions in fewer steps and with much higher returns compared to the KIX-A variant.
-It suggests that the agents learn higher-order relations and apply these learned relations and interactions with objects to solve problems.
+It suggests that the agents learn higher-order relations and apply these learned relations in interactions with objects to solve problems.
 Meta-level Guided Interactions Aid Agents
@@
 We evaluate the distribution of environmental returns for each agent across four tasks.
+Figure 5: Evaluation return profiles of agents performing tasks show that the agents benefit from meta-level guided interactions.
 For each evaluation task, trained agents are run for 12K episodes by parallel workers, and their environmental rewards are tracked.
-Figure 5: Evaluation return profiles of agents performing various tasks show that the proposed agents benefit from meta-level guided interactions.
 To understand the effects of meta-guided interactions and ensure a fair comparison between agents, regardless of success rate differences, we examine the top $k$ returns with $k=100$.
@@
 In this environment, the agent’s goal is to pick up a blue ball hidden in a two-room maze.
-The rooms are separated by a door that is obstructed by a ball and locked.
+The rooms are separated by a locked door that is obstructed by a ball.
 The keys are hidden inside boxes.
 Agents employ previously acquired interaction policies from the base task in the ObstructedMaze-Full environment while only learning the meta-policies.
-During evaluation, agents use the learned meta-policy, but the interaction policies are from the ObstructedMaze-Full environment itself.
-Figure 6 shows the meta training returns and evaluation return profiles, similar to the main environment, for both agent variants.
+In evaluation, agents use the learned meta-policy, but the interaction policies are from the ObstructedMaze-Full environment.
+Figure 6 shows evaluation return profiles for the base task, similar to the main environment, for both agent variants.
 It is evident that agents have gained the ability to exploit interaction policies from the previous environment, and the second variant, which employs a separate reachability policy, appears to have benefited more.
 Conclusion
-In this work, we proposed a metacognitive framework that promotes generalist behaviors through three main contributions.
-First, generalist behaviors necessitate abstraction and composition; however, current models lack natural abstractions and compositions that can preserve both object abstractions and relational knowledge structures.
-In our model, type-space graphs are used to represent and ground natural abstractions and compositions while also preserving relational knowledge between objects.
-Second, rather than acting directly in the environment, interactions with objects are critical in performing tasks.
+Artificial intelligence systems must develop knowledge-centric metacognitive abilities in the pursuit of abstract reasoning, knowledge learning, adaptive behavior and autonomy.
+The proposed metacognitive approach promotes these capabilities through three core contributions.
+First, knowledge learning and adaptation necessitate abstraction and composition while preserving relational knowledge structures.
+In our model, type graphs are used to represent and ground natural abstractions while preserving relational knowledge among objects.
+Second, rather than acting directly in the environment, interactions with objects are essential in task execution.
 To accomplish this, we introduced an intermediary interaction level.
-The agent’s meta-level interprets abstract relational knowledge, learns a meta-policy using type-space graphs, and makes recommendations on which objects to interact with and how to interact with them.
-In this sense, our model functions as an interpretation-based interaction goal-generating machine.
-Third, interactions employing interaction policies enable the necessary loose coupling required for generalization by acting as intermediaries.
-We demonstrated that metacognitive knowledge-guided interactions aid agents in generalization.
-Lastly, Our framework is modular and adaptable, enabling the integration of knowledge structures and control variables at both the meta and interaction levels.
-This framework is promising, particularly for domains where higher-order knowledge and interaction compositions can be exploited, such as robotics, autonomous and cognitive systems.
+At the meta level, the agent interprets abstract relational knowledge, learns a meta-policy over type graphs, and generates recommendations regarding which objects to interact with and how.
+In this sense, the proposed model functions as an interpretation-based interaction goal generating machine.
+Third, meta level guided interactions executed through interaction policies provide the loose coupling necessary for generalization by acting as intermediaries.
+The metacognitive approach facilitates knowledge learning of policies, knowledge guided interactions aid generalization and transferable interaction policies.
+Lastly, the proposed framework is modular and adaptable, enabling the integration of knowledge structures and control variables at both the meta and interaction levels.
+It is particularly promising for physical-world domains where abstract relational knowledge and compositional interactions can be exploited, such as robotics, autonomous, and cognitive systems.